### v1

# Zero-Shot Hyperspectral Pansharpening Using Hysteresis-Based Tuning for Spectral Quality Control

Abstract
Hyperspectral pansharpening has received much attention in recent years due to technological and methodological advances that open the door to new application scenarios. However, research on this topic is only now gaining momentum. The most popular methods are still borrowed from the more mature field of multispectral pansharpening and often overlook the unique challenges posed by hyperspectral data fusion, such as i) the very large number of bands, ii) the overwhelming noise in selected spectral ranges, iii) the significant spectral mismatch between panchromatic and hyperspectral components, iv) a typically high resolution ratio. Imprecise data modeling especially affects spectral fidelity. Even state-of-the-art methods perform well in certain spectral ranges and much worse in others, failing to ensure consistent quality across all bands, with the risk of generating unreliable results. Here, we propose a hyperspectral pansharpening method that explicitly addresses this problem and ensures uniform spectral quality. To this end, a single lightweight neural network is used, with weights that adapt on the fly to each band. During fine-tuning, the spatial loss is turned on and off to ensure a fast convergence of the spectral loss to the desired level, according to a hysteresis-like dynamic. Furthermore, the spatial loss itself is appropriately redefined to account for nonlinear dependencies between panchromatic and spectral bands. Overall, the proposed method is fully unsupervised, with no prior training on external data, flexible, and low-complexity. Experiments on a recently published benchmarking toolbox show that it ensures excellent sharpening quality, competitive with the state-of-the-art, consistently across all bands. The software code and the full set of results are shared online on https://github.com/giu-guarino/rho-PNN.
Index Terms:
Hyperspectral images, image fusion, pansharpening, spectral unmixing, zero-shot learning, self-supervised learning.
I Introduction
Hyperspectral (HS) remote sensing imagery is a key resource for monitoring and managing such diverse environments as forests [1, 2], agricultural areas [3, 4], oceans [5, 6], coastal areas [7, 8], inland waters and wetlands [9, 10], snow and ice [11, 12], atmosphere [13, 14], mineral resources [15, 16] [17], urban areas [18, 19]. Due to the technological constraints of the sensors and the very high number of spectral bands acquired, HS images have a rather low spatial resolution. A cost-effective solution to overcome this limitation is to equip the flight system with an additional camera to acquire, in synchrony with the HS image, a single high-resolution panchromatic (PAN) band. Then, using a fusion process known as pansharpening the PAN-HS pair is merged into a single high-resolution HS datacube. This data fusion technique enhances spatial resolution while preserving spectral consistency. Algorithms designed to improve the spatial resolution of remotely sensed images have already been shown to significantly boost performance in downstream tasks such as classification [20] and land use/land cover mapping [21]. It is therefore reasonable to expect that enhanced HS data could offer substantial benefits in applications where HS information is crucial but often underexploited due to limited spatial resolution, for example, in air pollution estimation [22] or water quality monitoring [23]. Examples of satellite-borne missions that feature the joint acquisition of PAN and HS images are PRISMA (PRecursore IperSpettrale della Missione Applicativa) by the Italian Space Agency (ASI), and EO-1/ALI (Earth Observing-1/Advanced Land Imaging) by the U.S. National Aeronautics and Space Administration (NASA). In both cases, the spatial resolution is 30m for the HS component and much better for the PAN image, 5m for PRISMA and 10m for EO-1/ALI. Needless to say, the PRISMA case is certainly more interesting, as it promises to provide high-resolution data, but also more challenging since it requires the pansharpening algorithm to handle a resolution ratio of 6 between the available HS data in input and the desired output.
Pansharpening is not a new topic. It has been the object of intense research in the last decades with reference to multispectral (MS) images, where only a limited number of spectral bands (from 4 to 8) are available in the visible to near-infrared spectrum. A large number of methods have been proposed for MS pansharpening [24, 25], with very different approaches such as component substitution [26, 27, 28, 29], multiresolution analysis [30, 31, 32], variational optimization [33, 34, 35], [36] and machine/deep learning [37, 38, 39, 40, 41]. Only recently the attention is shifting towards the HS case, as testified by the growing number of research papers (see [42] for a review) and by open challenges such as [43].
Clearly, MS and HS pansharpening share a common formulation which may suggest that PAN-MS fusion methods can be readily extended to operate also in the PAN-HS case. Indeed, early methods for HS pansharpening [44] were simple adaptations from the MS case, based on Bayesian models [45, 46, 47, 48], matrix factorization [49, 50, 51], variational optimization [52], component substitution and multiresolution analysis [53] (see [54] for a detailed review). However, this approach is overly simplistic, since the HS case has numerous peculiar features that make it particularly challenging:
  • PAN and MS images span comparable spectral ranges, and the PAN bandwidth covers most MS bands. The HS image, instead, has a much larger spectral extension and most of the HS bands are not covered by the PAN. Therefore, their spatial structure cannot be reliably predicted from it.
  • HS bands are much narrower than MS ones, in the order of 10nm. This is why, to harvest more energy, larger ground cells are necessary. Even so, the acquired data are rather noisy, with an extremely variable quality, and some groups of bands, in spectral ranges where the sunlight is weaker, must be discarded altogether.
  • Sometimes, the PAN-MS resolution ratio is rather large (e.g., six for PRISMA), hence the pansharpening algorithm is called upon to recover a large amount of missing information.
  • HS images are huge, hence the computational complexity is a serious issue, whatever the optimization strategy adopted. This is especially troublesome for supervised machine learning methods that require large training sets covering the whole data space.
In latest years, a few model-based methods specifically designed for HS pansharpening were proposed, relying on guided filtering [55], variational optimization [56, 57], saliency-based component substitution [58]. However, the deep learning (DL) revolution has by now spread to the field of remote sensing [59] and, since the seminal 2016 work of Masi et al. [37], deep learning is a de facto standard for pansharpening. Focusing on the HS case, in [60] and [61] dedicated convolutional neural networks (CNNs) were designed to strengthen the spectral prediction capability. In [62] a dual-attention residual network was proposed, with a deep HS image prior module for HS upsampling. In [63, 64] the problem of scale and resolution variability was tackled by means of a CNN with arbitrary-scale attention modules [64]. An overcomplete residual network that learns high-level features using constrained receptive fields was proposed in [65]. Multibranch network architectures were also explored in several recent works [66, 67, 68]. Other methods adopted classical pansharpening paradigms and used DL modules to estimate the model parameters [69] or to process the extracted features [70]. All these methods rely on supervised learning and obviate the absence of a real ground truth (GT) by training on reduced-resolution simulated images. The performance is indeed very good on reduced-resolution test data but drops significantly when moving to the real target of the process, the full-resolution images. In other words, models trained in the subsampled domain do not generalize well to the full-resolution domain. Recently, to overcome these problems, a paradigm shift towards full-resolution training is taking place for both MS [71, 72, 73] and HS [74, 75, 76] pansharpening, with novel learning strategies and suitably defined loss functions. Focusing on the HS case, the RE-RANet method proposed in [74] leverages the ratio image as latent variable and derives spatial and spectral information of the fused image from the full-resolution PAN and HS components. The R-PNN model [75] performs pansharpening sequentially for each individual band using an unsupervised target-adaptive solution. A divide-and-conquer strategy is used by PCA-Z-PNN, where the HS stack is split in two coherent sub-stacks which are transformed by principal component analysis and pansharpened individually by the MS-oriented Z-PNN [72] model. A common trait of these three methods is the use of two-term losses that suitably balance spectral and spatial quality.
Aiming to provide an objective overview of recent progress in the field and to support further research, we have recently developed a benchmarking toolbox [^1] https://github.com/matciotola/hyperspectral_pansharpening_toolbox for HS pansharpening [42]. A large number of state-of-the-art (SotA) methods, both model-based and data-driven, have been reimplemented using a common development environment and evaluated under uniform conditions on a large and representative set of publicly available images. In [42] we performed a comparative analysis of the state of the art. However, the toolbox can be used to perform many more experiments and investigate specific aspects of pansharpening, such as the strengths and weaknesses of different approaches. In particular, here we focus on the issue of non-uniform quality along the spectral dimension, a problem largely neglected in the literature. Indeed, we observed that nearly all methods perform well in some spectral intervals and much worse in others, providing unreliable results in some spectral bands. Of course, spectral richness is a key distinctive feature of HS images, but it becomes meaningless in the absence of adequate guarantees on the accuracy of the observed data. Only under these conditions, these images can be smoothly used by the end user. It should be noted that a major cause of this undesirable behavior is the lack of a quality assessment protocol specifically tailored to the hyperspectral case. The widely used performance metrics are simple adaptations of those designed for the very different MS case and, as such, they fail to capture subtle but important differences.
Motivated by the above considerations, in this work we propose a novel DL-based HS pansharpening method which is able to keep the spectral distortion under control in each single band, so as to ensure uniform (and top) quality along the whole spectral range. To this end, we build upon the recently proposed [75] Rolling Pansharpening Neural Network (Rolling PNN or R-PNN) method whose structure perfectly fits our need. R-PNN performs a sequential band-wise pansharpening using a single lightweight neural network which is adapted on the fly to the target band, with no prior training but for the first band. This architecture allows us to address individually each band’s needs. First, the desired level of spectral distortion is estimated based on the band features. Then a band-adaptive optimization schedule is enacted to reach this target level, where the spatial loss is carefully balanced with the spectral loss and switched on and off according to a hysteresis-like schedule in order to avoid undesired local minima. In addition, the spatial loss itself is re-defined to account for inverted PAN-HS correlations that can occur in specific spectral ranges. Experimental results on the benchmarking toolbox [42] prove the effectiveness of the proposed solution, which is competitive with the SotA on real full-resolution images in terms of global indexes and much superior in terms of spectral quality consistency.
In summary, the proposed method inherits the appealing properties of R-PNN and adds further strengths:
  • a.
    Fully unsupervised and zero-shot, with no need for prior training, no cross-dataset generalization issues.
  • b.
    Highly flexible: the number of spectral bands does not need to be fixed.
  • c.
    Uniform distortion along the whole spectral range, with target level defined by the user through a single and easily interpretable global parameter.
  • d.
    SotA performance on real full-resolution images, especially in terms of spectral quality.
  • e.
    Scalable complexity, which allows to trade-off computation for quality based on application needs.
In the rest of the paper, Section II provides some necessary background concepts, Section III describes the proposed solution, Section IV presents and comments comparative experimental results, and Section V draws conclusions and outlines future perspectives.
II Background
The solution proposed in this paper moves from our recent Rolling PNN method [75], which we briefly summarize here, highlighting its distinctive features. Then we analyze the issue of spectral fidelity for current state of the art methods.
II-A Rolling PNN
Figure 1: R-PNN flowchart. Each module (a) inherits the initial weights, $\phi_{b-1}$, from the previous module, fine-tunes them on the current band, and passes the updated version, $\phi_{b}$, to the next module, as depicted in (b). The $b$-th band is pansharpened with its own tuned parameters $\phi_{b}$. $N_{b}$ is the number of tuning iterations assigned to band $b$.
TABLE I: Main symbols
Symbol Dimensions Meaning
$w,h$ Scalars HS width and height, respectively
$W,H$ Scalars PAN width and height, respectively
$R$ Scalar Resolution ratio ($R=H/h=W/w$)
$B$ Scalar # of HS bands
$\mathbf{H}$ $[w{\times}h{\times}B]$ Original HS image
$\mathbf{P}$ $[W{\times}H]$ Original PAN image
$\widetilde{\mathbf{H}}$ $[W{\times}H{\times}B]$ $R{\times}R$ upscaled version of $\mathbf{H}$
$\widehat{\mathbf{H}}$ $[W{\times}H{\times}B]$ Pansharpened version of $\mathbf{H}$
$\mathbf{D}$ $[W{\times}H{\times}B]$ Estimated details from $\mathbf{P}$ and $\widetilde{\mathbf{H}}$
[…]
${\cal L}_{\lambda}$, ${\cal L}_{S}$, ${\cal L}$ Scalars Spectral, spatial and overall loss functions
Compared to the more familiar case of pansharpening of MS images, in the HS case the generalization of deep learning models is usually more challenging for several reasons: less training data, richer spectral information, low correlation between PAN and spectral bands outside the visible spectrum, larger resolution ratios, and variability of the number of bands due to several technical reasons. Starting from the above observations and inspired by [39, 72], we proposed R-PNN [75], a CNN-based band-wise pansharpening solution where model parameters are not pre-trained but generated on the fly on the same target image at inference time. Unlike previous works, suited to the MS case, in R-PNN the PAN is fused with only one band at a time, rather than all of them, and the tuned model is passed from one band to another for targeted adaption. The overall band-wise pansharpening process is summarized in Fig. 1, with the main symbols used here and in the reminder of the paper described in Tab. I.
In mathematical terms, the network of Fig. 1 (a) performs the mapping $\widehat{\mathbf{H}}_{b}=f(\mathbf{H}_{b},\mathbf{P};\phi)$, where $\phi$ is a set of trainable parameters. For each band, $b$ (except band 1 starting from pretrained parameters), a tuning process starts from the initial parameters, $\phi^{(0)}=\phi_{b-1}$, inherited from the previous band, and runs for $N_{b}$ iterations to get the tuned parameters $\phi_{b}=\phi^{(N_{b})}$. $N_{b}$ is proportional (up to a bounding value) to the spectral distance between bands $b$ and $b-1$. In fact, HS datasets often present spectral gaps due to discarded bands, and longer training helps making up for reduced band similarity.
Each tuning iteration is carried out on the same components $\mathbf{H}_{b}$ and $\mathbf{P}$ to be fused, thanks to an unsupervised multitask loss which balances spectral and spatial consistencies of the fused image $\widehat{\mathbf{H}}_{b}$:
${\cal L}={\cal L}_{\lambda}\left(\widehat{\mathbf{H}}_{b},\mathbf{H}_{b}\right)+\beta{\cal L}_{S}\left(\widehat{\mathbf{H}}_{b},\mathbf{P}\right),$ (1)
with $\beta$ a balancing hyperparameter. The spectral consistency is given by
${\cal L}_{\lambda}\left(\widehat{\mathbf{H}}_{b},\mathbf{H}_{b}\right)=\|\widehat{\mathbf{H}}_{b}^{(\downarrow)}-\mathbf{H}_{b}\|_{1}$ (2)
where $\widehat{\mathbf{H}}_{b}^{(\downarrow)}$ is the downscaled (lowpass filtered and decimated) version of $\widehat{\mathbf{H}}_{b}$. The spatial consistency, instead, is given by the average local correlation between $\widehat{\mathbf{H}}_{b}$ and $\mathbf{P}$. Given the correlation coefficient $\rho(s)$, computed on a small $\sigma{\times}\sigma$ neighborhood of pixel $s$, the loss is defined as
${\cal L}_{S}\left(\widehat{\mathbf{H}}_{b},\mathbf{P}\right)=\langle[\rho^{\rm max}(s)-\rho(s)]_{+}\rangle$ (3)
where $\langle x\rangle$ indicates spatial average, $[x]_{+}=\max(0,x)$, and $\rho^{\rm max}(s)$ is a rough estimate of the expected local correlation, computed on reduced resolution data. Therefore, this spatial loss aims at maximizing $\rho(s)$, but only as long as it does not exceed a reasonable target value, $\rho^{\rm max}(s)$.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 2: Reprojection spectral error curves, $E_{b}$, for the PRISMA images Cagliari (a) and Udine (b), and corresponding normalized curves $e_{b}$ (c) and (d).
II-B Analysis of spectral distortion
For all methods of the benchmarking toolbox [42], and for two different PRISMA test images, we computed the band-wise reprojection error [^2] Although this quantity is often referred to as spectral distortion, here we prefer to recall the presence of a downscaling step using the term reprojection.
$E_{b}=\|\widehat{\mathbf{\mathbf{H}}}_{b}^{(\downarrow)}-\mathbf{H}_{b}\|_{1}$ (4)
obtained by comparing the downscaled (smoothed and decimated) version of the pansharpened image with the input HS image. This is the same distance used in R-PNN (2) but other metrics provide similar results. In Fig. 2 (a)-(b) we plot results only for the methods characterized by the best full-resolution spectral quality.
The error is clearly band-dependent. Almost all methods follow a common pattern, certainly related also to the different levels and properties of noise in different bands. Even the simple EXP interpolator, which is a 23-tap polynomial approximation of the ideal interpolator, follows the same dynamics. The average level of each curve provides information about the global distortion incurred by the method (remember that we are neglecting spatial quality in this analysis) but the detailed behavior of the curves shows that numbers vary significantly from band to band, and that the average distortion may be a poor indicator of the actual distortion observed in certain frequency ranges. Under this point of view, a uniform level of distortion would be certainly an appreciable feature.
To simplify the analysis of results let us consider a normalized version of the same data. Since the EXP interpolator works exclusively on $\mathbf{H}$, without adding any high-frequency spatial detail coming from the PAN, it can be regarded as a method-independent “spectral reference”. Therefore, we use its reprojection error $E^{\rm EXP}_{b}$ to relativize the profiles of Fig. 2 by defining a normalized error $e_{b}$ as
$e_{b}=E_{b}\,/\,E^{\rm EXP}_{b}$ (5)
The resulting curves are plot in Fig. 2 (c)-(d). Spectral error profiles are now easier to analyze. As “virtuous” example we can mention TV which, for both test images, keeps the normalized reprojection error low (about 60% of EXP) and almost constant over the whole spectral range. In contrast, other methods show significant imbalances between different spectral ranges. This may be partly due to spatial detail injection rules that do not properly account for the spectral-spatial interdependence. For example, several methods achieve very low spectral distortions in the visible range (first $\sim$30 bands) with a sharp performance impairment beyond the near infrared limit, possibly due to the inappropriate injection of spatial details in a range where the correlation between PAN and HS bands decreases. On the opposite side we mention the case of PCA-Z-PNN, which follows a divide-and-conquer approach where the HS bands are split in two groups and separately processed. This split is clearly reflected in the almost piece-wise uniform $e_{b}$ curve (see Fig. 2 (d)), with a relatively large distortion in the visible range and a much better behavior beyond that range.
Overall, this preliminary analysis motivates us to work on a pansharpening method that adapts to the different characteristics exhibited by HS bands in different spectral ranges and provides nearly uniform quality across the spectrum. R-PNN, with its band-wise optimization procedure, is an excellent candidate to achieve this goal.
III Proposed Method
III-A Correlation-based Spatial Loss for HS Images
In our proposal, we keep the two-component loss formulation (1) previously adopted [73] in unsupervised pansharpening and use the same spectral component of [75], based on the $\ell_{1}$-norm (2), which shows good convergence properties and outperforms other norms, such as $\ell_{2}$ or ERGAS, in terms of perceptual quality. In any case, the exact form of the spectral loss is not really critical as it is always obtained by comparing the pansharpened product $\widehat{\mathbf{\mathbf{H}}}$ with a clear reference, the interpolated HS image $\widetilde{\mathbf{\mathbf{H}}}$ or (after decimation) the low-resolution HS itself. We focus instead on the spatial loss term, which is more elusive as it is intimately related to the problem of assessing spatial quality without a reference. In R-PNN we used the correlation-based loss (3) originally proposed in [72] for MS pansharpening. The idea is to push the pansharpened band to have a high local correlation with the PAN, so as to exhibit the same spatial layout. The correlation coefficient measures how well one image can be linearly predicted from another. A coefficient of $|\rho|=1$ indicates perfect predictability, while $\rho=0$ indicates no correlation. Ideally, pansharpened bands should closely match the spatial structure of the PAN image, resulting in high correlation. However, to preserve spectral information, this correlation cannot be unitary. In fact, it is expected to vary spatially and among different bands, depending on the scene’s characteristics. For this reason, to prevent the injection of alien details from the PAN into the spectral band, the correlation itself is capped by its estimate, $\rho^{\rm max}(s)$, computed on reduced resolution data.
This correlation-based loss proved extremely effective for MS pansharpening, ensuring results characterized by very high spatial fidelity [72, 73]. Its extension to the HS case, however, is more controversial. In fact, the basic assumption is that $\rho(s)$ is generally high and certainly positive. While this is more than reasonable for the MS case, it becomes largely questionable in the HS case because many spectral bands fall outside the visible range (i.e., the PAN range) and exhibit a lower correlation with it, and sometimes the so-called inversion phenomenon [77] with regions having strong but negative correlation. An algorithm that disregards these physical principles may produce pansharpened results with visible artifacts, such as unnatural color variations and exaggerated contrasts. Furthermore, during the target adaptation phase, this inconsistency could lead to a higher conflictual behavior between the spectral and spatial loss functions, ultimately hindering the overall reconstruction quality.
To clarify this point, we carried out a correlation analysis on a sample PRISMA image (Kansas). Lacking the ideal pansharpened datacube, we estimate the local correlation on reduced resolution data, using the canonical resolution downgrade of the PAN to align it with the available HS datacube. We select two sample spectral bands which represent opposite behaviors, one in the visible spectrum, thus highly correlated with the PAN, the other outside this range and with much lower correlation. In Fig. 3 (top row) we show the downgraded PAN $\mathbf{P}^{(\downarrow)}$, the two spectral bands, $\mathbf{H}_{v}$ (from the visible spectrum) and $\mathbf{H}_{o}$ (outside), together with the corresponding correlation maps, $\rho_{v}=\rho_{\mathbf{P}^{(\downarrow)}\mathbf{H}_{v}}$ and $\rho_{o}=\rho_{\mathbf{P}^{(\downarrow)}\mathbf{H}_{o}}$, respectively. The rightmost image is a quantized representation of $\rho_{o}$ used to simplify its interpretation. On the bottom row of the figure, the same items are shown for a meaningful close-up of the image (green box).
PAN: $\mathbf{P}^{(\downarrow)}$   $\mathbf{H}_{v}$ $\rho_{v}$, $\langle\rho_{v}\rangle=0.95$   $\mathbf{H}_{o}$ $\rho_{o}$, $\langle\rho_{o}\rangle=0.33$ $\rho_{o}$ (quantized)
{Overpic}Refer to caption  {Overpic}Refer to caption {Overpic}Refer to caption  {Overpic}Refer to caption {Overpic}Refer to caption {Overpic}Refer to caption
Refer to caption   Refer to caption Refer to caption   Refer to caption Refer to caption Refer to caption
$-1.0<\rho_{o}\leq-0.6$     $-0.6<\rho_{o}\leq-0.2$     $-0.2<\rho_{o}\leq 0.2$     $0.2<\rho_{o}\leq 0.6$     $0.6<\rho_{o}\leq 1.0$
Figure 3: PAN-HS local correlation analysis. First row, from left to right: PAN image shifted at the HS spatial resolution ($\mathbf{P}^{(\downarrow)}$), two sample HS bands paired with the corresponding map of local correlation with the PAN, from visible ($\mathbf{H}_{v},\rho_{v}$) and NIR-SWIR ($\mathbf{H}_{o},\rho_{o}$) spectral ranges, and (rightmost) the quantized version of the correlation map $\rho_{o}$. Second row: close-up on a meaningful detail (highlighted with a green box).
A few observations are in order. The sample band from the visible spectrum shows a very high correlation with the PAN, 0.95 on the average and never negative. This is a common behavior in the visible spectrum, more pronounced in the green range and a little less towards the coastal and red/infrared limits. In this situation, maximizing the local correlation, without any upper bound, makes perfect sense. Outside the visible spectrum, however, the correlation field has a much higher dynamics, and our sample band has an average correlation of just 0.33 with the PAN. Such a low value, however, does not signal lack of dependence. The spatial layouts keep being mostly the same, as easily appreciated in the close-up image, but several regions exhibit a strong negative correlation. This is the case, for example, of the circularly shaped elements clearly visible in the close-up, which are very dark in the PAN, and very bright in $\mathbf{H}_{o}$. In addition, intermediate correlation values, between -0.6 and 0.6, occur at the boundaries between positively and negatively correlated regions just because of the finite size of the estimation window (6$\times$6 in our experiment). This latter problem impacts also on the estimation of the bounding map, $\rho^{\rm max}$, further reducing its effective resolution and making it scarcely reliable.
Based on the above observations, we decided to define a spatial loss term that maximizes the absolute value of the correlation with the PAN, so as to promote structural similarity irrespective of the correlation polarity.
${\cal L}_{S}\left(\widehat{\mathbf{H}}_{b},\mathbf{P}\right)=\left\langle 1-|\rho(s)|\right\rangle$ (6)
Note that we have also removed the bounding map $\rho^{\rm max}$ due to its limited reliability. However, we will show that a suitable hysteresis-based optimization schedule will prevent the injection of inappropriate PAN details in the pansharpened image.
Armed with our new loss, comprising the original spectral term (2) and the new spatial term (6), we can now pursue our original goal, that is, devising a pansharpening algorithm that provides a uniform quality across the whole spectrum. More precisely, considering that the various bands experience very different noise intensities, we aim at achieving a uniform relative quality, measured by the error normalized with respect to the EXP ideal interpolator (5).
III-B Network Architecture
As said before, we rely on rolling pansharpening [75] as it allows us to operate band-wise. With this strategy, only one low-resolution interpolated spectral band is given in input, together with the PAN, and only one high-resolution spectral band is provided in output. Therefore, any pansharpening network can be adapted to the purpose by setting appropriately the depth of the input and output layers. The network must be trained from scratch for the first band and then refined for each new band by target-adaptive tuning. Of course, for an HS image, this latter process must be repeated many times, which calls for the intrinsic requisite of a lightweight network.
We have implemented and tested several low-complexity solutions from the MS pansharpening literature and verified experimentally that network structure and capacity have a minor impact on performance provided some minimal features are guaranteed: i) no less than three convolutional layers; ii) a global receptive field of a reasonable size; iii) use of a global skip connection (residual net). Given the above indications, we decided to simply keep the network used in R-PNN, a single-band residual version of PNN [39], which meets all these criteria. Tab. II summarizes the network layers and connections.
TABLE II: Network layers and topology.
Layer Size Kernel Outputs
Input $W{\times}H{\times}2$ $(\widetilde{\mathbf{\mathbf{H}}}_{b},\mathbf{P})$
Conv+ReLU $W{\times}H{\times}48$ $7{\times}7$
Conv+ReLU $W{\times}H{\times}32$ $7{\times}7$
Conv $W{\times}H{\times}1$ $5{\times}5$ $\mathbf{D}_{b}$
Output (add gate) $W{\times}H{\times}1$ $\widetilde{\mathbf{\mathbf{H}}}_{b}+\mathbf{D}_{b}$
III-C Hysteresis-based Training Procedure
Figure 4: Loss trajectories in the $({\cal L}_{\lambda/E},{\cal L}_{S})$ plane when tuning models with various choices of the $\alpha$ and $\beta$ parameters. Tuning starts from random initial weights and proceeds for a large number (1000) of iterations. Sample bands from the visible range (top), and outside it (bottom).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Loss trajectories in the $({\cal L}_{\lambda/E},{\cal L}_{S})$ plane when tuning models with various choices of the $\alpha$ and $\beta$ parameters. Tuning starts from random initial weights and proceeds for a large number (1000) of iterations. In each chart, the curves refer to two different bands, B1 from the visible spectrum (violet) and B2 outside it (orange), and two different tuning schedules, the proposed histeresis-like strategy (solid lines) and flat tuning (dashed lines).
Once fixed the network architecture, we can focus on optimizing network training and fine tuning. To this end we can play with several hyperparameters, first of all the learning rate $\alpha$, the weight $\beta$ that balances the spectral and spatial terms in the loss, and the number of iterations $N_{b}$. To gain insight into the problem, we carried out some preliminary experiments using various $(\alpha,\beta)$ pairs and studying their impact on the two loss components, ${\cal L}_{S}$ and ${\cal L}_{\lambda}$ (these losses depend also on the band, $b$, but we avoid a further subscript for the sake of clarity). To avoid initialization issues, we trained the network from scratch and used a very large number of iterations, $N_{b}=1000$. In Fig. 4, for two different bands, one from the visible range (top) and one outside it (bottom), we plot the evolution of the loss components in the $({\cal L}_{\lambda/E},{\cal L}_{S}$) plane as a function of the number of iterations. For the reasons explained before, for the spectral loss we consider the normalized value ${\cal L}_{\lambda/E}={\cal L}_{\lambda}/{\cal L}^{\rm EXP}_{\lambda}$. Each curve starts from the top-right point and evolves over time towards lower values of the losses.
For the “visible” band everything works quite well. At the end of the process, ${\cal L}_{S}$ is always very low, which is to be expected given the very high correlation with the PAN, and also the spectral quality is quite good, with ${\cal L}_{\lambda/E}$ going largely below 1 for some $(\alpha,\beta)$ pairs, thus improving significantly over the reference value provided by the EXP method. For the “outside” band the story is completely different. ${\cal L}_{S}$ never approaches 0, and this was expected considering the much lower correlation of this band with the PAN. However, also ${\cal L}_{\lambda}$ remains rather large and only in some cases it approaches the reference ${\cal L}^{\rm EXP}_{\lambda}$ level. The loss trajectories are especially informative. In all cases, ${\cal L}_{S}$ improves regularly as the process goes on but this seems to affect negatively ${\cal L}_{\lambda}$ which, after an initial improvement, impairs markedly and never recovers. Contrary to what happens for the visible band, ${\cal L}_{\lambda/E}$ never goes below 1.
This behavior gave us the idea to use a more sophisticated training schedule, where the balance between spectral and spatial loss components is adjusted dynamically during the process. In particular, we decided to set $\beta=0$ at the beginning and fine-tune the network based only on the spectral loss, so as to bring it down to a desired level, say ${\cal L}_{\lambda/E}=\gamma_{{}_{\rm LOW}}$. Then we turn the ${\cal L}_{S}$ loss component back on, but always keeping the spectral quality under control. In fact, as soon as ${\cal L}_{\lambda/E}$ exceeds a second higher threshold, $\gamma_{{}_{\rm HIGH}}$, the spatial loss component is turned off again, and ${\cal L}_{\lambda/E}$ is brought again to the lower level. Our intuition was that this optimization cycle could bring the curve to a new point in the trajectory, with the same spectral loss as before but a lower spatial loss. Fig. 5 shows that this is actually the case. In this figure, we plot the same curves as before but organized differently. Each chart now refers to a different $(\alpha,\beta)$ pair and compares the $({\cal L}_{\lambda/E},{\cal L}_{S})$ trajectories obtained with flat training and with the proposed on-off schedule. For the visible band, the proposed schedule does not really improve over the baseline (the mechanism could be even switched off). On the contrary, for the outside band the gain is striking. The desired level of spectral distortion is readily achieved with all $(\alpha,\beta)$ pairs. Then ${\cal L}_{S}$ is gradually reduced while ${\cal L}_{\lambda}$ keeps oscillating in the desired range. The final value of ${\cal L}_{S}$ is somewhat larger than with flat training, but this is quite reasonable for an “outside” band, intrinsically less correlated with the PAN, and does not correspond to a worse spatial quality. A further interesting property emerging from this analysis is the remarkable stability of the trajectories, almost identical to one another as the hyperparameters change, which suggests a good robustness with respect to changing sources, namely, a good generalization ability. Finally, note that most of the final iterations are used to gain only minor improvements in ${\cal L}_{S}$, and tuning could be safely stopped much earlier. This last observation allowed us to define a computationally efficient stopping rule. For the current band, $b$, we count both the total number of iterations $N_{b}$ and the number of iterations with active spatial loss ($\beta>0$), $N_{b|S}$. Tuning stops when either one exceeds a predefined limit, whose setting is discussed in next Section. Note that also the value of $\beta$ to be used when the spatial loss is active is not fixed once and for all but only after a suitable warm-up. A large value, $\beta_{0}$, is used initially, but this is repeatedly halved if it causes an excessive increase in spectral loss, down to a minimum of $\beta_{0}/8$. This mechanism relieves us from the burden of selecting a correct value for $\beta$ which on the other hand can change from band to band. Even so, a few parameters remain to be set, a problem that will be analyzed in some detail in the following Section.
Figure 6: Close-up (bands 55 to 59) of the evolution of spectral (a) and spatial (b) loss terms during the hysteresis-based tuning with model transfer.
Fig. 6 illustrates an example of tuning with hysteresis for a few bands of a sample image. Parts (a) and (b) show the evolution of the (normalized) spectral and, respectively, spatial loss components as the tuning proceeds. Red lines highlight ON phases, where both loss terms are active. In this example we have fixed $\gamma_{{}_{\rm LOW}}=0.72$ and $\gamma_{{}_{\rm HIGH}}=0.8$. Several interesting observations can be made: i) most of the times, the model optimized for the previous band works very well on the new one, e.g. 55$\to$56 or 58$\to$59. In some cases, e.g. 56$\to$57, the transfer is less effective which impacts significantly on quality and calls for a tuning overhead to bring the spectral loss within the hysteresis limits. This latter case, however, occurs rarely, typically when going from the last band of a correlation block to the first band of another one. ii) different average levels of spatial loss are observed for different bands, reflecting their different relationship with the PAN. iii) boosting sharpness (${\cal L}_{S}$ ON) causes an increase of the spectral loss ${\cal L}_{\lambda}$ and the need for multiple hysteresis cycles to keep it within the desired limits.
Given the oscillatory behavior of both losses, and some overshooting due to momentum, the last model provided by the fine-tuning procedure is not necessarily the best performing one. Therefore, in our implementation, we keep memory of all of them and output the model with the lowest spatial loss among those that satisfy the constraints on the spectral loss.
IV Experimental Analysis
In this Section, we perform the comparative performance assessment of the proposed method. First, we set some key parameters of the method based on preliminary experiments on the validation set of the benchmarking toolbox [42]. Then, after discussing the assessment methodology, we carry out an objective performance analysis on reduced-resolution and full-resolution test data of the same toolbox, and finally perform visual analysis of results.
IV-A Parameter Setting
Figure 7: Loss points in the $({\cal L}_{\lambda/E},{\cal L}_{S})$-plane for various bands after thorough hysteresis-based tuning as a function of the $\gamma_{{}_{\rm HIGH}}$ and $\gamma_{{}_{\rm LOW}}$ parameters.
IV-A1 Setting the Target for Spectral Loss
a key question for our approach is: to what extent can the spectral distortion be reduced without causing appreciable spatial distortion? To answer this question, we carried out a set of experiments on five bands {B1, …, B5} of the validation dataset, sampled uniformly from the whole spectral range. For each band, we trained the model from scratch (no cross-band model transfer) with our hysteresis-based schedule, for virtually infinite iterations, using always $\alpha=10^{-5}$ and $\beta=0.5$, but varying $\gamma_{{}_{\rm HIGH}}$ in $\{0.40,0.45,\ldots,0.90\}$, with $\gamma_{{}_{\rm LOW}}=0.9\gamma_{{}_{\rm HIGH}}$.
In Fig. 7 we show the final loss values achieved as points in the $({\cal L}_{\lambda/E},{\cal L}_{S})$ plane. For all bands we observe a similar behavior. The lowest spatial loss is obtained for $\gamma_{{}_{\rm HIGH}}=0.9$ and differs from band to band depending on its noise level and especially its similarity with the PAN. For bands in the visible spectrum, e.g. B1, with very high local correlation with the PAN, ${\cal L}_{S}$ goes almost to 0 (0.03), while it is much higher for other bands, e.g. B2, from the near-IR range, with ${\cal L}_{S}$ close to 0.2. In all cases, the spatial loss increases as $\gamma_{{}_{\rm HIGH}}$ decreases, as expected, but just slowly at the beginning, and a steep increase is observed only when ${\cal L}_{\lambda/E}$ falls below 0.5, that is, ${\cal L}_{\lambda}$ is 50% of the reference ${\cal L}^{\rm EXP}_{\lambda}$ value. Based on this consistent behavior, we decided to set $\gamma_{{}_{\rm HIGH}}=0.65$ and $\gamma_{{}_{\rm LOW}}=0.59$ in the upcoming experiments.
IV-A2 Limiting the Number of Iterations
Figure 8: R-PNN normalized reprojection error over four PRISMA images.
besides ensuring high and consistent quality, we want to keep complexity under control by limiting the number of tuning iterations. However, as also shown by Fig. 6, the previous-band model is not always a good starting point, in which case many iterations are required to bring the loss down to acceptable values. This appears clearly in Fig. 8 which shows the normalized reprojection error of R-PNN over four PRISMA images. The error spikes over the first band of a correlation block, then stabilizes at lower values in subsequent bands due to the high inter-band similarity. Based on this observation, we decided to stop tuning when either one of two conditions holds:
$\{N_{b|S}=N_{b|S}^{\rm max}\}\;\;\mbox{OR}\;\;\{N_{b}=N^{\rm max}_{b}\},$ (7)
Most of the times, the spectral loss reaches quickly the desired value, and most iterations involve also the spatial loss. In these cases, the active constraint is on $N_{b|S}$, and we set the maximum to a fixed value for all bands, $N_{b|S}^{\rm max}=N_{S}^{\rm max}$. More rarely, when subsequent bands are weakly correlated, reducing the spectral loss is more difficult, and a limit on the total number of iterations $N^{\rm max}_{b}$ becomes necessary to prevent an inordinate number of them. This latter limit, however, must be carefully set to make sure that a reasonable number of iterations involve also the spatial loss. Therefore, we decided to make this limit depend on inter-band correlation, that is
$N^{\rm max}_{b}=N_{0}+\Delta N(1-c_{b})$ (8)
with $c_{b}$ the correlation coefficient between bands $b$ and $b-1$. Therefore, the maximum number of iterations is $N_{0}\geq N_{S}^{\rm max}$ in case of perfect correlation ($c_{b}=1$) and $N_{0}+\Delta N$ in case of orthogonal bands ($c_{b}=0$). With this rule, we can easily set an overall bound to the total number of iterations $N$ as follows
$N=\sum_{b}N_{b}\leq N^{\rm max}=\sum_{b}N^{\rm max}_{b}=(1+\eta)BN_{0}$ (9)
being $\eta$ a suitable overhead multiplier to be set by the user, and $B$ the number of bands. Therefore, one can set $N^{\rm max}$ in advance based on complexity issues, and $N_{0}$ and $\eta$ accordingly. Then, $\Delta N$ will be given by the (easy to proof) formula
$\Delta N=\frac{\eta BN_{0}}{B-\sum_{b}c_{b}}$ (10)
Fig. 9 shows an example of how iterations distribute over the various bands of a PRISMA image. Here we set $N_{S}^{\rm max}=20,N_{0}=50$ and $\eta=8$. $N_{b|S}$ is shown in yellow and $N_{b}$ in brown, with its upper bound $N^{\rm max}_{b}$ in light gray. Although a maximum of $(1+\eta)N_{0}=450$ iterations per band is allowed on average, only 47 are used on the average. Most of the times, tuning stops because $N_{b|S}$ reaches its maximum, $N_{S}^{\rm max}=20$. The total number of iterations rarely exceeds $N_{0}=50$ and only for a group of bands, on the boundary between the visible and NIR ranges, the $N_{b}=N_{b}^{\rm max}$ exit condition is activated. The entire pansharpening procedure of the proposed method is summarized in Algorithm 1.
Figure 9: Distribution of tuning iterations for a test PRISMA image. On average, 47 iterations/band are used, almost always 20 of them with spatial loss ON.
Tab. III summarizes the most relevant hyperparameters of the proposed method, their selected values, and their meaning.
Algorithm 1 Pseudo-code $\rho$-PNN
Hyperparameters =$=\{\gamma_{high},\gamma_{low},\beta_{0},\eta,N_{0},N_{S}^{max},\epsilon,\alpha\}$
$B\leftarrow size(H,3)$triangleright # of bands
$c\leftarrow correlations(H)$triangleright vector $c_{b}=corrcoef(H_{b},H_{b-1})$; $c_{0}=0$
$\phi\leftarrow weightInit(net)$triangleright random initialized net parameters
for $b\leftarrow 1,B$ do
    $N_{b}^{max}\leftarrow iterationUpperBound(c,N_{0},B,\eta)$triangleright eqs 8 and 10
    $\tilde{H}_{b}\leftarrow expInterpolator(H_{b})$
    $\mathcal{L}_{\lambda}^{EXP}\leftarrow l_{1}\text{-}norm(H_{b},subsample(\tilde{H}_{b}))$triangleright reprojection error on $\tilde{H}_{b}$
    $N_{b},N_{b|s}\leftarrow 0$triangleright counters initializations
    $\beta\leftarrow warmUp(H_{b},P,net,\phi,\beta_{0},\epsilon,\alpha)$triangleright warmup to fix $\beta$
    $full\leftarrow{{}^{\prime}OFF^{\prime}}$triangleright full loss (’ON’) vs spectral-only loss (’OFF’)
    while  $\{N_{b|S}<N_{S}^{max}\}\;\;\mbox{OR}\;\;\{N_{b}<N^{\rm max}_{b}$}  do
        $\phi,\mathcal{L}_{\lambda},\mathcal{L}_{S}\leftarrow tuningStep(H_{b},P,net,\phi,\alpha,\beta,full)$
        $N_{b}\leftarrow N_{b}+1$
        if {full} then
           $N_{b|S}\leftarrow N_{b|S}+1$
           if $\mathcal{L}_{\lambda}/\mathcal{L}_{\lambda}^{EXP}>\gamma_{high}$then
               $full\leftarrow^{\prime}OFF^{\prime}$triangleright swith-off
           else if $\mathcal{L}_{\lambda}/\mathcal{L}_{\lambda}^{EXP}<\gamma_{low}$then
               $full\leftarrow{{}^{\prime}ON^{\prime}}$triangleright swith-on
           end if
        end if
    end while
    $\hat{H}_{b}\leftarrow predict(H_{b},P,net,\phi)$triangleright inference on band $b$
end for
return $\hat{H}$
TABLE III: Main hyperparameters of the proposed method.
Param Value Meaning
$\gamma_{{}_{\rm HIGH}}$ 0.65 Superior limit fixed for ${\cal L}_{\lambda/E}$
$\gamma_{{}_{\rm LOW}}$ 0.59 Inferior limit for ${\cal L}_{\lambda/E}$
$\beta_{0}$ 2 Starting $\beta$, weight of the spatial loss term
$\epsilon$ 0.007 Tolerated growth factor of ${\cal L}_{\lambda}$ for $\beta$ adaptation
$N_{S}^{\rm max}$ 20 Target # of tuning iterations with active spatial loss
$N_{0}$ 80 Max # of iterations for bands with perfect correlation with the preceding one
$\eta$ 30 Overhead factor to globally constrain extra iterations
IV-B Assessment Methodology
To assess the quality of the proposed solution we take advantage of the Benchmarking toolbox [42] which provides a publicly available set of rich and diverse PAN-HS PRISMA images, a large set of SotA comparative methods, and several performance evaluation tools based on the most credited assessment indexes. PRISMA was chosen for this evaluation since it provides a real panchromatic image acquired through a dedicated sensor operating in the 400 - 700 nm spectral range. In addition to the PAN sensor, PRISMA is equipped with two hyperspectral instruments: a VNIR sensor covering wavelengths from 400 to 1010 nm with 66 bands, and a SWIR sensor spanning 920 to 2505 nm with 173 bands, resulting in a total of 239 spectral channels. The dataset released by [42] consists of 16 PRISMA images, 12 designated for training and validation, and 4 for testing. To enhance data reliability, noisy spectral bands are removed, reducing the dataset to 159 clean and informative bands for benchmarking purposes. Additionally, to further investigate the generalizability of the proposed method, we considered an additional test dataset in which the PAN image was simulated using the RGB bands of the Copernicus Sentinel-2 mission (665,560 and 490 nm with 10m spatial resolution) and subsequently co-registered with the PRISMA HS data. Tab. IV gathers the comparative methods which can be roughly grouped in four categories: component substitution (CS), multiresolution analysis (MRA), model-based optimization (MBO), and deep learning (DL). The interested Reader is referred to [42] for details on these methods.
TABLE IV: HS pansharpening methods.
Name Ref Summary
EXP Approximation of the ideal interpolator
Component Substitution (CS) methods
GSA [26] Gram-Schmidt adaptive component substitution
BT-H [29] Brovey transform with haze correction
BDSD-PC [78] Band-dependent spatial detail injection with physical constraint
PRACS [28] Partial replacement adaptive CS
Multiresolution Analysis (MRA) methods
MTF-GLP-FS [79] Modulation Transfer Function (MTF)-matched Generalized Laplacian Pyramid (MTF-GLP) with fusion rule at full scale
MTF-GLP-HPM [80] MTF-GLP with high pass modulation
MTF-GLP-HPM-R [81] MTF-GLP-HPM with regression-based spectral matching
AWLP [30] Additive wavelet luminance proportional
MF [32] Nonlinear decomposition with morphological filters
Model-Based Optimization (MBO) methods
HySURE [46] Bayesian estimation with vector total variation prior
SR-D [82] Sparse representations-based detail injection
TV [33] Total variation-based pansharpening
Deep Learning (DL) methods
HyperPNN [60] 7-layer net with spectral encoder-decoder structure
HSpeNet [61] Advanced version of HyperPNN, with deeper architecture and Spectral Angle Mapper (SAM) loss term
DHP-DARN [62] Deep residual channel-spatial attention net with Deep Image Prior (DIP) for HS resizing
DIP-HyperKite [65] Overcomplete net with DIP for HS resizing
HyperDSNet [83] Spectral attention-based detail injection from deep-shallow features
R-PNN (unsup.) [75] Band-wise pansharpening using modified Z-PNN [72] with tuning propagation
PCA-Z-PNN (unsup.) [76] Z-PNN model with PCA-based input reduction
The assessment follows the well-established Wald’s protocol [84] which considers two contexts: a reduced resolution framework, where suitably degraded (in resolution) datasets are used for assessment with GT (where the ground truth is given by the original full resolution HS image), and a full resolution framework for assessment without GT but using the original data as input. Accordingly, the quality indexes, gathered in Tab. V, are split in two groups for reduced resolution (RR) and full resolution (FR) assessment.
As well known, the assessment of spatial quality in the full resolution context (hence, without a GT) is still an open and challenging problem. Therefore, to obtain a more comprehensive evaluation of performance, besides the $D_{S}$ index used in the toolbox [42] we provide also the correlation-based index $D_{\rho}$ proposed in [85]. Even so, no numerical index or combination of indexes provides a fully reliable measure of spatial quality, and we will resort to visual inspection of sample results to get further indications on spatial quality.
TABLE V: HS pansharpening quality assessment indexes.
RR assessment
ERGAS Erreur Relative Globale Adimensionnelle de Synthése [86]
SAM Spectral Angle Mapper [87]
$Q2^{n}$ Multiband extension [88] of Universal Image Quality Index [89]
FR assessment
$D_{\lambda}$ Khan’s spectral distortion index [90, 91]
$D_{S}$ Spatial distortion index [90, 92]
$D_{\rho}$ Correlation-based spatial distortion index [85]
TABLE VI: PAN-MS PRISMA benchmark datasets for testing.
Date Location Land covers
05/09/22 Cagliari, Italy Roofs, Streets, Crops, Hills, Water
24/08/23 Udine, Italy Roofs, Streets, Crops, Water
08/09/23 Ford County, Kansas Streets, Crops
20/11/23 Macuspana, Tabasco Streets, Crops
RR test image size: 600$\times$600 (from a 3600$\times$3600 real FR image)
Real FR test image size: 1200$\times$1200
Resolution ratio $R=6$
Bands $B=159/239$
Resolution: 30 m (RR) or 5 m (FR)
Experiments are carried out on the test datasets of the Toolbox, summarized in Tab. VI, comprising four acquisitions/scenes. From each original PRISMA PAN-HS pair (PAN/HS size: 3600$\times$3600/600$\times$600) two test datasets were created: a FR one (PAN/HS size: 1200$\times$1200/200$\times$200) obtained by cropping; a RR one (PAN/HS/GT size 600$\times$600/100$\times$100/600$\times$600) obtained by resolution downgrading of the whole PAN-HS pair, with the original FR HS playing the role of GT. Regarding the additional test dataset with the Sentinel-2 simulated PAN, we selected a Sentinel-2 acquisition covering the same area as the PRISMA Kansas dataset (in Tab. VI), acquired with a one-day temporal shift, and co-registered it accordingly. To preserve the original PRISMA ratio of 6, the hyperspectral data were then downscaled by a factor of 2, resulting in a spatial resolution of 60 meters (PAN/HS size: 600$\times$600/100$\times$100).
IV-C Reduced-resolution Quality Assessment
The numerical results obtained for the four RR datasets are summarized in Tab. VII. For each index, for each dataset and on average (Avg.), the top 5 and the worst 5 methods are highlighted in green and red, respectively. The best one is in bold green. Before analyzing the results, focusing on DL methods, it is important to recall that supervised models are optimized for the RR case while unsupervised ones (R-PNN, PCA-Z-PNN, and proposed) are optimized for FR real data. Therefore, it is not surprising that several supervised techniques (e.g., HyperPNN, HSpeNet and especially Hyper-DSNet) are on top. However, despite its unsupervised approach, the proposed solution also performs quite well. More specifically, it presents a good spectral behavior (SAM is the most appropriate index in this respect) and also a very good capacity to preserve spatial structures ($Q2^{n}$ is an index oriented to assess spatial correlation and contrast). It looks less competitive in terms of ERGAS, a band-wise normalized variant of the root mean square error, but still close to the top 5.
TABLE VII: Results at Reduced Resolution. Best score is in bold green, Top-5 in green, and Worst-5 in red.
[…]
  Cagliari Udine Ford Tabasco Avg.   Cagliari Udine Ford Tabasco Avg.   Cagliari Udine Ford Tabasco Avg.
(Ideal)   0 0 0 0 0   0 0 0 0 0   1 1 1 1 1
EXP   1.7716 3.9587 1.6921 4.3650 2.9468   2.3073 4.6288 2.8528 6.3524 4.0353   0.5971 0.6453 0.7984 0.5935 0.6586
GSA   0.9389 2.4243 1.0008 2.8065 1.7926   1.8191 3.5730 2.2570 5.7455 3.3486   0.8739 0.8110 0.9201 0.7955 0.8501
BT-H   1.2784 3.4898 1.2462 3.1015 2.2790   2.3857 4.6059 2.5563 6.3532 3.9753   0.8180 0.7166 0.8941 0.7837 0.8031
BDSD-PC   1.1664 2.7627 1.2637 3.9316 2.2811   1.9439 5.0116 2.6646 7.8908 4.3777   0.8524 0.7904 0.9048 0.7144 0.8155
PRACS   1.1803 3.4688 1.2773 3.8578 2.4461   1.9742 4.2221 2.4571 6.2161 3.7174   0.8166 0.7144 0.8804 0.6778 0.7723
MTF-GLP-FS   0.9239 2.3660 0.9808 2.7356 1.7516   1.8123 3.4496 2.2297 5.7208 3.3031   0.8792 0.8218 0.9238 0.8026 0.8568
MTF-GLP-HPM   0.9292 4.0148 1.2231 3.0609 2.3070   1.8555 5.1824 2.6120 6.4104 4.0151   0.8833 0.6698 0.9000 0.7890 0.8105
MTF-GLP-HPM-R   0.9114 2.4661 0.9743 2.6967 1.7621   1.8121 3.3953 2.2124 6.0986 3.3796   0.8800 0.8134 0.9239 0.8023 0.8549
AWLP   1.1846 4.6926 1.3424 3.8592 2.7697   2.3323 7.9280 2.9252 7.7668 5.2381   0.8516 0.6290 0.8832 0.7496 0.7783
MF   1.1392 4.0155 1.5506 3.3944 2.5249   1.9213 4.5587 2.7173 6.2918 3.8723   0.8638 0.7263 0.8804 0.7807 0.8128
HySURE   1.4664 4.3476 2.0749 5.1119 3.2502   2.9264 6.6950 4.4153 9.0468 5.7709   0.7943 0.5761 0.7475 0.4951 0.6532
SR-D   1.8476 4.2618 1.8825 4.8011 3.1982   2.4242 6.1149 3.0551 7.9927 4.8967   0.5734 0.5846 0.7313 0.5440 0.6083
TV   1.2624 2.8864 1.3163 3.1771 2.1605   2.2212 4.0400 2.5705 5.6672 3.6247   0.8291 0.7833 0.8911 0.7692 0.8182
HyperPNN   1.2245 2.2237 1.1864 3.3557 1.9976   2.8925 4.3058 2.6461 7.0296 4.2185   0.8777 0.8539 0.9247 0.8229 0.8698
HSpeNet   1.0354 2.0527 1.0674 2.7964 1.7379   2.0644 3.2157 2.2838 5.8462 3.3525   0.8674 0.8588 0.9136 0.8245 0.8660
DHP-DARN   1.9137 4.4280 1.7389 4.7617 3.2106   3.4673 5.1740 3.3833 9.1288 5.2883   0.8489 0.8234 0.9190 0.7934 0.8462
DIP-HyperKite   1.3510 2.0206 1.0995 2.8092 1.8201   3.6543 5.8268 2.7089 9.1932 5.3458   0.8419 0.8497 0.9205 0.8201 0.8581
Hyper-DSNet   0.9725 1.8198 0.9820 2.4237 1.5495   1.9162 2.9638 2.1906 7.4854 3.6390   0.8833 0.8767 0.9303 0.8426 0.8832
R-PNN   0.9884 2.2256 1.0581 2.6704 1.7356   1.9720 3.5728 2.3860 5.6797 3.4026   0.8712 0.8225 0.9172 0.8112 0.8555
PCA-Z-PNN   1.0824 2.9062 1.5575 2.7813 2.0819   2.1124 4.6125 3.1154 5.5297 3.8425   0.8507 0.7798 0.8647 0.8116 0.8267
$\rho$-PNN   1.0109 2.3268 1.0205 2.7308 1.7722   1.9625 3.5609 2.2439 5.3396 3.2767   0.8782 0.8279 0.9246 0.8177 0.8621
TABLE VIII: Results at Full Resolution. Best score is in bold green, Top-5 in green, and Worst-5 in red
. Method   $D_{\lambda}$   $D_{S}$   $D_{\rho}$   Cagliari Udine Ford Tabasco Avg.   Cagliari Udine Ford Tabasco Avg.   Cagliari Udine Ford Tabasco Avg. (Ideal)   0 0 0 0 0   0 0 0 0 0   0 0 0 0 0 EXP   0.0126 0.0077 0.0054 0.0106 0.0091   0.1306 0.1273 0.0704 0.1947 0.1308   0.7973 0.8839 0.7963 0.7972 0.8187 GSA   0.0227 0.0134 0.0078 0.0122 0.0140   0.0056 0.0074 0.0077 0.0184 0.0098   0.0593 0.6009 0.2075 0.0791 0.2367 BT-H   0.0633 0.0936 0.0195 0.0288 0.0513   0.0000 0.0000 0.0000 0.0002 0.0001   0.0579 0.0954 0.0544 0.0676 0.0688 BDSD-PC   0.0439 0.0256 0.0220 0.0487 0.0351   0.0024 0.0000 0.0001 0.0003 0.0007   0.1280 0.3878 0.1726 0.1192 0.2019 PRACS   0.0118 0.0155 0.0047 0.0088 0.0102   0.0083 0.0163 0.0114 0.0245 0.0151   0.1775 0.3252 0.3409 0.1903 0.2585 MTF-GLP-FS   0.0101 0.0054 0.0033 0.0065 0.0063   0.0214 0.0277 0.0174 0.0322 0.0247   0.0618 0.6000 0.2123 0.0831 0.2393 MTF-GLP-HPM   0.0147 0.0087 0.0067 0.0093 0.0098   0.0215 0.0284 0.0179 0.0389 0.0267   0.0479 0.1022 0.0438 0.0538 0.0619 MTF-GLP-HPM-R   0.0100 0.0057 0.0033 0.0074 0.0066   0.0217 0.0285 0.0176 0.0374 0.0263   0.0669 0.6050 0.2164 0.0950 0.2458 AWLP   0.0132 0.0097 0.0055 0.0092 0.0094   0.0251 0.0318 0.0206 0.0407 0.0295   0.0695 0.1778 0.0754 0.0933 0.1040 MF   0.0655 0.0392 0.0390 0.0331 0.0442   0.0389 0.0269 0.0250 0.0664 0.0393   0.0476 0.0904 0.0391 0.0585 0.0589 HySURE   0.1126 0.0618 0.0510 0.1158 0.0853   0.0022 0.0011 0.0016 0.0022 0.0018   0.1414 0.2276 0.0990 0.2688 0.1842 SR-D   0.0128 0.0087 0.0055 0.0115 0.0096   0.1316 0.1292 0.0720 0.1981 0.1327   0.8234 0.9061 0.8531 0.8560 0.8596 TV   0.0042 0.0028 0.0021 0.0036 0.0032   0.0586 0.0524 0.0358 0.0813 0.0570   0.2590 0.3294 0.2791 0.2001 0.2669 HyperPNN   0.0415 0.0139 0.0256 0.0248 0.0264   0.0036 0.0063 0.0040 0.0067 0.0051   0.1310 0.5080 0.2419 0.2625 0.2859 HSpeNet   0.0174 0.0084 0.0240 0.0154 0.0163   0.0101 0.0160 0.0141 0.0236 0.0159   0.1439 0.5338 0.2348 0.2558 0.2921 DHP-DARN   0.0832 0.0615 0.0467 0.0513 0.0607   0.0030 0.0053 0.0031 0.0130 0.0061   0.1494 0.4713 0.2106 0.2315 0.2657 DIP-HyperKite   0.0208 0.0125 0.0148 0.0146 0.0157   0.0042 0.0050 0.0043 0.0059 0.0048   0.2952 0.5270 0.2217 0.2765 0.3301 Hyper-DSNet   0.0119 0.0050 0.0082 0.0068 0.0080   0.0200 0.0246 0.0195 0.0346 0.0247   0.1957 0.5336 0.2276 0.3353 0.3230 R-PNN   0.0131 0.0091 0.0063 0.0074 0.0090   0.0168 0.0144 0.0129 0.0341 0.0195   0.0511 0.1366 0.0532 0.0504 0.0728 PCA-Z-PNN   0.0129 0.0041 0.0046 0.0080 0.0074   0.0152 0.0116 0.0088 0.0257 0.0153   0.0959 0.6828 0.7935 0.0792 0.4128 $\rho$-PNN   0.0043 0.0027 0.0018 0.0043 0.0033   0.0215 0.0168 0.0173 0.0380 0.0234   0.0790 0.1302 0.0743 0.0532 0.0842
TABLE IX: Results at Full Resolution on Sentinel-2 / PRISMA dataset. Best score is in bold green, Top-5 in green, and Worst-5 in red
. Method   $D_{\lambda}$ $D_{S}$ $D_{\rho}$ (Ideal)   0 0 0 EXP   0.0039 0.0985 0.7919 GSA   0.0068 0.0002 0.2833 BT-H   0.0338 0.0000 0.0298 BDSD-PC   0.0095 0.0011 0.1236 PRACS   0.0047 0.0023 0.2629 MTF-GLP-FS   0.0019 0.0134 0.2820 MTF-GLP-HPM   0.0039 0.0140 0.0384 MTF-GLP-HPM-R   0.0019 0.0139 0.2861 AWLP   0.0034 0.0154 0.0626 MF   0.0278 0.0340 0.0439 HySURE   0.0553 0.0068 0.1323 SR-D   0.0043 0.1015 0.8624 TV   0.0018 0.0536 0.2117 HyperPNN   0.0824 0.0027 0.1967 HSpeNet   0.0920 0.0033 0.0799 DHP-DARN   0.1136 0.0035 0.3505 DIP-HyperKite   0.0044 0.0032 0.1657 Hyper-DSNet   0.0284 0.0121 0.1894 R-PNN   0.0032 0.0079 0.0516 PCA-Z-PNN   0.0013 0.0067 0.8200 $\rho$-PNN   0.0013 0.0071 0.0521
IV-D Full-resolution quality assessment
FR numerical results are gathered in Tab. VIII using the same highlighting scheme of Tab. VII. The first five columns give the spectral quality index, $D_{\lambda}$, individually on the four test images and averaged on all of them. The next two groups of columns, instead, measure spatial quality by means of two alternative indices, $D_{S}$ and $D_{\rho}$. In terms of spectral quality, the proposed method provides top scores, on par with TV, and much better than all other methods, including its ancestor R-PNN. This was our aim from the beginning, together with that of uniform quality across the whole spectrum, not catched by this index.
$D_{\lambda}$ is a solid indicator of quality, as it is computed against a well defined and unambiguous reference. Measuring spatial quality is much more challenging, which is why we consider two alternative indices. In fact, neither of them is without flaws. For example, $D_{S}$ quantifies the distance between the PAN image and the optimal linear combination of the pansharpened spectral bands that best approximates it [92]. This is physically motivated by the fact that a high-resolution hyperspectral (HS) sensor would capture narrowband signals that collectively contribute to the broadband response of the PAN sensor. Therefore, an ideal set of sharpened HS bands should be able to reconstruct the PAN image through an appropriate linear combination. However, if one band (among hundreds) is an exact copy of the PAN, then $D_{S}=0$, no matter what the other bands look like. This happens for several entries in Tab. VIII which, far from signaling perfect pansharpening, correspond typically to degenerate solutions. While such cases are rarely seen with MS images, they occur more easily with HS images. Similar warnings apply to the second indicator. In fact, $D_{\rho}$ is defined as the average of the local (calculated on a $\sigma{\times}\sigma$ domain) correlation coefficient between PAN and the pansharpened HS bands [85]. From a physical perspective, this metric evaluates the extent to which a $\sigma{\times}\sigma$ patch from any band of $\widehat{\mathbf{\mathbf{H}}}$ can be linearly predicted from the corresponding patch in the PAN image. Consequently, $D_{\rho}$ directly reflects the degree of spatial alignment between the pansharpened product $\widehat{\mathbf{\mathbf{H}}}$ and the PAN image $P$ at fine spatial scales. Therefore, it vanishes only when there is perfect correlation between all bands and the PAN. However, this is not desirable for bands that have no spectral overlap with the PAN and may differ significantly from it. Although both indices have their limitations, we believe that their combined use offers a more comprehensive assessment of spatial consistency. Keeping this in mind, we comment on the numerical results with some caution and await confirmation from the visual analysis to draw firm conclusions. In particular, the proposed method performs quite well in terms of spatial quality indices, especially when considering $D_{\rho}$. Together with the excellent results of $D_{\lambda}$, this speaks of a reliable and well-balanced method. This does not apply to TV, however, whose performance drops dramatically, suggesting questionable spatial quality. On the contrary, MF is characterized by excellent spatial quality indices but a very bad $D_{\lambda}$. Both these methods, and many others, seem unable to achieve a good balance between spectral and spatial fidelity. Overall, the proposed method compares quite favorably with all DL-based references. Only R-PNN provides a better $D_{\rho}$ on average, but this is also due to the perfect alignment between the metric and the spatial loss of R-PNN, a property that does not hold with the new spatial loss used in the proposal. The numerical results for the Sentinel-2/PRISMA dataset are reported in Table IX. In this experiment, it is evident that none of the implemented methods, except for the proposed one, achieves a top-5 ranking in both spectral quality ($D_{\lambda}$) and spatial quality ($D_{S}$ and/or $D_{\rho}$).
IV-E Visual results
The numerical results of Tab. VII provide indications on RR performances that are clear and reliable, although only proxies of the actual performances in the FR space. These latter are quantified directly in Tab. VIII. However, while $D_{\lambda}$ captures spectral quality unambiguously, $D_{S}$ and $D_{\rho}$ provide more controversial and often contrasting indications on spatial quality. Therefore, we rely on the visual inspection of some sample results to gain a more solid insight on this latter point.
In Fig. 10 and Fig. 11 we show results for two closeups from the FR Cagliari and the FR Tabasco datasets. Following [42], we show two different spectral triplets of bands for false-RGB composition, from the visible and NIR-SWIR spectral ranges. Given the large number of reference methods, we limit attention to the top-5 $D_{S}$ and top-5 $D_{\rho}$ methods (based on averages from Tab. VIII) plus TV, top-1 $D_{\lambda}$, for a total of 10 methods including the proposed one. [^3] The full set of results is available on the web repository at https://github.com/giu-guarino/rho-PNN.. Figure 12 presents the results for the Sentinel-2/PRISMA dataset. As in Figures 10 and 11, we include 10 methods following the same selection criteria, with the exception of replacing TV with PCA-Z-PNN (the second-best performer in terms of $D_{\lambda}$ after the proposed method). Together with the images, we also report the spectral quality ($D_{\lambda}$) and spatial quality ($D_{S}/D_{\rho}$) scores computed over all bands of the full scene. Additionally, Figure 13 presents the visual results of the top-5 methods in terms of $Q^{2n}$ on the RR Udine dataset, along with their corresponding error maps. As in previous cases, we also report the evaluation metrics ($ERGAS$, $SAM$, and $Q^{2n}$) computed over all spectral bands and the full scene.
PAN HS BT-H BDSD-PC HySURE DIP-HyperKite
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HyperPNN TV MF MTF-GLP-HPM R-PNN $\rho$-PNN
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Pansharpening results on the FR Cagliari dataset (240$\times$360 close-up): PAN image (top left) with paired lower resolution HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab. VIII).
PAN HS BT-H BDSD-PC HySURE DIP-HyperKite
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HyperPNN TV MF MTF-GLP-HPM R-PNN $\rho$-PNN
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Pansharpening results on the FR Tabasco dataset (240$\times$360 close-up): PAN image (top left) with paired lower resolution HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab. VIII).
PAN HS GSA BT-H BDSD-PC PRACS
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HyperPNN MTF-GLP-HPM MF R-PNN PCA-Z-PNN $\rho$-PNN
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Pansharpening results on the Sentinel-2 / PRISMA dataset (240$\times$360 close-up): Simulated PAN image obtained from Sentinel-2 RGB bands (top left) with paired lower resolution PRISMA HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab. VIII).
GT HyperPNN HSpeNet DIP-HyperKite Hyper-DSNet $\rho$-PNN
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Pansharpening results on the RR Udine dataset (120$\times$180 close-up) and error maps (Top 5 $Q2^{n}$). RGB (a) and False color (b) results are displayed. For each result the corresponding scores “ERGAS, SAM, $Q2^{n}$” are reported.
From the visual inspection of images we can make the following observations:
  • Results with low $D_{\rho}$ are generally sharper, resembling closely the spatial structures of the PAN. This is true for both PRISMA FR scenes and Sentinel-2/PRISMA scene and in both the visible and NIR-SWIR spectral ranges. In particular, this is the case of BT-H, MF, MTF-GLP-HPM, R-PNN and proposed.
  • Results with low $D_{S}$ reproduce the PAN spatial structures faithfully in the visible spectrum, but seem to fail outside that range. This is the case of BDSD-PC, HySURE and DIP-HyperKite form PRISMA FR scene, whereas GSA, BDSD-PC, and PRACS are the cases for Sentinel-2/PRISMA scene. BT-H seems an exception, characterized also by a low $D_{\rho}$, however it presents a large spectral distortion.
  • For PRISMA FR scenes, the top-$D_{\lambda}$ method, TV, whose spectral quality level is nearly the same as the proposed one, does not provide effective sharpening, providing blurred images.
For the RR evaluation framework, we selected image crops from the Udine area (see Fig. 13), displaying results for the visible bands in the top row and the NIR–SWIR bands in the bottom row. The left image shows the ground truth (i.e., the original HS image), followed by the pansharpening results of the five best-performing methods according to the Q2n metric, and finally, the output of our proposed method. The visual quality of the results produced by the top-5 algorithms is generally comparable with that of our method, confirming the trends observed in the quantitative results reported in Tab. VII. However, a closer inspection reveals that our method tends to preserve spectral fidelity more effectively. In contrast, the output of HyperPNN exhibits noticeable color artifacts. Although some of the other methods provide slightly sharper spatial details, they also introduce visible spectral aberrations, particularly in homogeneous regions such as the sea.
Based on the above observations, we single out three methods, MTF-GLP-HPM, R-PNN and our proposal, that appear to be both robust and effective, performing uniformly well on all scenes and all displayed spectral bands in terms of both spectral and spatial quality. Samples of these methods are shown close to each other but the differences are so small that further examination seems like a futile exercise that could lead to claims related only to the choice of scene/bands/clips. On the other hand, all three methods have very similar spatial quality indicators. However, we point out that the proposed method has a much smaller average spectral distortion (0.0033, in Tab.VIII and 0.0013 in Tab.IX) than these competitors (0.0098 and 0.0090, respectively, in Tab.VIII, and 0.0039 and 0.0032, respectively in Tab.IX). Furthermore, only the proposed method ensures uniform quality across all bands, a feature not captured by the global indicators.
In summary, the proposed method demonstrates more consistent and robust performance compared to current SotA. It effectively preserves spectral fidelity without compromising spatial quality, and unlike other deep learning-based approaches, it does not require any training data for pansharpening inference. Furthermore, as confirmed by an additional test involving a PAN image acquired from a different sensor on a different date, the method exhibits a strong degree of generalizability. On the downside, under the current hyperparameter configuration, the approach entails a relatively high computational cost compared to both model-based and several deep learning-based techniques. However, by relaxing the spectral quality constraints (via $\gamma_{\text{high}}$ and $\gamma_{\text{low}}$), it is possible to accelerate the algorithm significantly.
V Conclusions
In this work we have proposed a new pansharpening method aimed at ensuring a uniformly high fidelity over the whole spectral range. This is an important feature for hyperspectral images, where most bands are acquired at spectral frequencies not covered by the PAN, and may differ significantly from it. To this end, we leverage the rolling PNN architecture where pansharpening is carried out band-wise using cross-band transfer learning and a suitable zero-shot unsupervised tuning. From this base, we improve with two key innovations: a new loss, better adapted to the characteristics of HS images, that takes into account the correlation inversion phenomena that often occur in bands outside the visible range; and a hysteresis-based tuning scheme where the balance of spatial and spectral loss components changes dynamically to overcome unwanted low-quality solutions.
Advantages
  • It works from scratch (Only PCA-Z-PNN does the same);
  • It works with an arbitrary number of bands (Only PCA-Z-PNN and R-PNN do the same);
  • It ensures a good generalization ability across diverse sensors;
  • It works at full resolution: No downgrade nor simulation operations are needed (same as R-PNN and PCA-Z-PNN);
  • It ensrures a controllable and uniform level of spectral control along all wavelengths.
Drawbacks
  • It is slower than pre-trained methods at inference time;
Extensive experimental assessment over a SotA benchmarking toolbox proves the proposed method to ensure excellent results both in terms of visual quality and according to the most credited quality indicators.
However, this latter point deserves certainly further research. To date, no spatial quality metric can be considered a fully reliable indicator of image quality at full-resolution. Lacking a solid unambiguous reference, numerical assessment of performance remains inconclusive. This is the most urgent research topic in the pansharpening context, in our opinion, as it is intimately linked to the definition of an effective loss for unsupervised learning.

### v2

@@
 I Introduction
-Hyperspectral (HS) remote sensing imagery is a key resource for monitoring and managing such diverse environments as forests [1, 2], agricultural areas [3, 4], oceans [5, 6], coastal areas [7, 8], inland waters and wetlands [9, 10], snow and ice [11, 12], atmosphere [13, 14], mineral resources [15, 16] [17], urban areas [18, 19].
+Hyperspectral (HS) remote sensing imagery is a key resource for monitoring and managing such diverse environments as forests [1, 2], agricultural areas [3, 4], oceans [5, 6], coastal areas [7, 8], inland waters and wetlands [9, 10], snow and ice [11, 12], atmosphere [13, 14], mineral resources [15, 16, 17], urban areas [18, 19].
 Due to the technological constraints of the sensors and the very high number of spectral bands acquired, HS images have a rather low spatial resolution.
@@
 Then, using a fusion process known as pansharpening the PAN-HS pair is merged into a single high-resolution HS datacube.
-This data fusion technique enhances spatial resolution while preserving spectral consistency.
-Algorithms designed to improve the spatial resolution of remotely sensed images have already been shown to significantly boost performance in downstream tasks such as classification [20] and land use/land cover mapping [21].
-It is therefore reasonable to expect that enhanced HS data could offer substantial benefits in applications where HS information is crucial but often underexploited due to limited spatial resolution, for example, in air pollution estimation [22] or water quality monitoring [23].
+This data fusion process improves spatial resolution while preserving spectral coherence, thereby boosting the performance of downstream tasks such as classification [20] and land use/cover mapping [21].
 Examples of satellite-borne missions that feature the joint acquisition of PAN and HS images are PRISMA (PRecursore IperSpettrale della Missione Applicativa) by the Italian Space Agency (ASI), and EO-1/ALI (Earth Observing-1/Advanced Land Imaging) by the U.S.
@@
 The idea is to push the pansharpened band to have a high local correlation with the PAN, so as to exhibit the same spatial layout.
-The correlation coefficient measures how well one image can be linearly predicted from another.
-A coefficient of $|\rho|=1$ indicates perfect predictability, while $\rho=0$ indicates no correlation.
-Ideally, pansharpened bands should closely match the spatial structure of the PAN image, resulting in high correlation.
-However, to preserve spectral information, this correlation cannot be unitary.
-In fact, it is expected to vary spatially and among different bands, depending on the scene’s characteristics.
-For this reason, to prevent the injection of alien details from the PAN into the spectral band, the correlation itself is capped by its estimate, $\rho^{\rm max}(s)$, computed on reduced resolution data.
+Indeed, the correlation coefficient measures how well one image can be linearly predicted from another.
+Since pansharpened bands and PAN share the same spatial structure, they can be expected to be strongly correlated.
+On the other hand, each band has unique spectral properties that must be carefully preserved.
+Therefore, to prevent the injection of alien details from the PAN into the spectral band, the correlation itself, $\rho$, is capped by its estimate, $\rho^{\rm max}(s)$, computed on reduced resolution data.
 This correlation-based loss proved extremely effective for MS pansharpening, ensuring results characterized by very high spatial fidelity [70, 71].
@@
 While this is more than reasonable for the MS case, it becomes largely questionable in the HS case because many spectral bands fall outside the visible range (i.e., the PAN range) and exhibit a lower correlation with it, and sometimes the so-called inversion phenomenon [75] with regions having strong but negative correlation.
-An algorithm that disregards these physical principles may produce pansharpened results with visible artifacts, such as unnatural color variations and exaggerated contrasts.
-Furthermore, during the target adaptation phase, this inconsistency could lead to a higher conflictual behavior between the spectral and spatial loss functions, ultimately hindering the overall reconstruction quality.
 To clarify this point, we carried out a correlation analysis on a sample PRISMA image (Kansas).
@@
 This latter problem impacts also on the estimation of the bounding map, $\rho^{\rm max}$, further reducing its effective resolution and making it scarcely reliable.
+An algorithm that disregards these phenomena may generate visible pansharpening artifacts, such as unnatural color variations and exaggerated contrasts.
+Moreover, conflicts may arise between the spectral and spatial terms of the loss, hindering the training or fine-tuning phases.
 Based on the above observations, we decided to define a spatial loss term that maximizes the absolute value of the correlation with the PAN, so as to promote structural similarity irrespective of the correlation polarity.
@@
 Note that we have also removed the bounding map $\rho^{\rm max}$ due to its limited reliability.
-However, we will show that a suitable hysteresis-based optimization schedule will prevent the injection of inappropriate PAN details in the pansharpened image.
+In fact, a suitable hysteresis-based optimization schedule will prevent the injection of inappropriate PAN details in the pansharpened image.
 Armed with our new loss, comprising the original spectral term (2) and the new spatial term (6), we can now pursue our original goal, that is, devising a pansharpening algorithm that provides a uniform quality across the whole spectrum.
@@
 Therefore, in our implementation, we keep memory of all of them and output the model with the lowest spatial loss among those that satisfy the constraints on the spectral loss.
+Algorithm 1 describes the proposed method more formally by means of a high-level pseudo-code.
+Input: $\mathbf{P}$, $\mathbf{H}$, net (network structure)
+Hyperparameters: $\gamma_{{}_{\rm HIGH}}$, $\gamma_{{}_{\rm LOW}}$, $\beta_{0}$, $\eta$, $N_{0}$, $N_{S}^{\rm max}$, $\epsilon$
+Output: $\widehat{\mathbf{\mathbf{H}}}$
+1:$B\leftarrow{\rm size}(\mathbf{H},3)$triangleright # of bands
+2:$\mathbf{c}\leftarrow{\rm correlations}(\mathbf{H})$triangleright $\mathbf{c}_{b}={\rm corrcoef}(\mathbf{H}_{b},\mathbf{H}_{b-1})$; $\mathbf{c}_{1}=0$
+3:$\phi\leftarrow{\rm weightInit}({\rm net})$triangleright net weights init.
+4:for $b\leftarrow 1:B$ do
+5:    $N_{b}^{\rm max}\leftarrow{\rm iterationUpperBound}(\mathbf{c},N_{0},B,\eta)$triangleright Eqs. (8), (10)
+6:    $\widetilde{\mathbf{H}}_{b}\leftarrow{\rm interpolator}(\mathbf{H}_{b})$
+7:    $\mathcal{L}_{\lambda}^{\rm EXP}\leftarrow{\rm getSpectralLoss}(\widetilde{\mathbf{\mathbf{H}}}_{b},\mathbf{H}_{b})$
+8:    $N_{b},N_{b|S}\leftarrow 0$triangleright counter init.
+9:    $\beta\leftarrow{\rm warmUp}(\mathbf{H}_{b},\mathbf{P},{\rm net},\phi,\beta_{0},\epsilon,\alpha)$triangleright $\beta$ warm-up
+10:    ${\rm ON}\leftarrow\mathbf{\rm false}$triangleright ${\cal L}_{S}$ flag init.
+11:    while  $\{N_{b|S}<N_{S}^{\rm max}\}\;\mbox{\bf and}\;\{N_{b}<N^{\rm max}_{b}$}  do $\triangleright$ Eq. (7)
+12:        $\phi,\mathcal{L}_{\lambda},\mathcal{L}_{S}\leftarrow{\rm optimizationStep}(\mathbf{H}_{b},\mathbf{P},{\rm net},\phi,\beta,{\rm ON})$
+13:        $N_{b}\leftarrow N_{b}+1$
+14:        if ON then
+15:           $N_{b|S}\leftarrow N_{b|S}+1$
+16:        end if
+17:        if ${\cal L}_{\lambda/E}>\gamma_{{}_{\rm HIGH}}$ then
+18:           ON $\leftarrow\mathbf{\rm false}$triangleright switch-off
+19:        else if ${\cal L}_{\lambda/E}<\gamma_{{}_{\rm LOW}}$ then
+20:           ON $\leftarrow\mathbf{\rm true}$triangleright switch-on
+21:        end if
+22:    end while
+23:    $\widehat{\mathbf{H}}_{b}\leftarrow{\rm predict}(\mathbf{H}_{b},\mathbf{P},{\rm net},\phi)$triangleright inference on band $b$
+24:end for
+25:return $\widehat{\mathbf{H}}$
+Algorithm 1 $\rho$-PNN pseudo-code.
 IV Experimental Analysis
@@
 Therefore, we decided to make this limit depend on inter-band correlation, that is
-$N^{\rm max}_{b}=N_{0}+\Delta N(1-c_{b})$ (8)
+$N^{\rm max}_{b}=N_{0}+\Delta N(1-\mathbf{c}_{b})$ (8)
 with $c_{b}$ the correlation coefficient between bands $b$ and $b-1$.
-Therefore, the maximum number of iterations is $N_{0}\geq N_{S}^{\rm max}$ in case of perfect correlation ($c_{b}=1$) and $N_{0}+\Delta N$ in case of orthogonal bands ($c_{b}=0$).
+Therefore, the maximum number of iterations is $N_{0}\geq N_{S}^{\rm max}$ in case of perfect correlation ($\mathbf{c}_{b}=1$) and $N_{0}+\Delta N$ in case of orthogonal bands ($\mathbf{c}_{b}=0$).
 With this rule, we can easily set an overall bound to the total number of iterations $N$ as follows
@@
 Then, $\Delta N$ will be given by the (easy to proof) formula
-$\Delta N=\frac{\eta BN_{0}}{B-\sum_{b}c_{b}}$ (10)
+$\Delta N=\frac{\eta BN_{0}}{B-\sum_{b}\mathbf{c}_{b}}$ (10)
 Fig.
@@
 The total number of iterations rarely exceeds $N_{0}=50$ and only for a group of bands, on the boundary between the visible and NIR ranges, the $N_{b}=N_{b}^{\rm max}$ exit condition is activated.
-The entire pansharpening procedure of the proposed method is summarized in Algorithm 1.
 Figure 9: Distribution of tuning iterations for a test PRISMA image. On average, 47 iterations/band are used, almost always 20 of them with spatial loss ON.
 Tab. III summarizes the most relevant hyperparameters of the proposed method, their selected values, and their meaning.
-Algorithm 1 Pseudo-code $\rho$-PNN
-Hyperparameters =$=\{\gamma_{high},\gamma_{low},\beta_{0},\eta,N_{0},N_{S}^{max},\epsilon,\alpha\}$
-$B\leftarrow size(H,3)$triangleright # of bands
-$c\leftarrow correlations(H)$triangleright vector $c_{b}=corrcoef(H_{b},H_{b-1})$; $c_{0}=0$
-$\phi\leftarrow weightInit(net)$triangleright random initialized net parameters
-for $b\leftarrow 1,B$ do
-    $N_{b}^{max}\leftarrow iterationUpperBound(c,N_{0},B,\eta)$triangleright eqs 8 and 10
-    $\tilde{H}_{b}\leftarrow expInterpolator(H_{b})$
-    $\mathcal{L}_{\lambda}^{EXP}\leftarrow l_{1}\text{-}norm(H_{b},subsample(\tilde{H}_{b}))$triangleright reprojection error on $\tilde{H}_{b}$
-    $N_{b},N_{b|s}\leftarrow 0$triangleright counters initializations
-    $\beta\leftarrow warmUp(H_{b},P,net,\phi,\beta_{0},\epsilon,\alpha)$triangleright warmup to fix $\beta$
-    $full\leftarrow{{}^{\prime}OFF^{\prime}}$triangleright full loss (’ON’) vs spectral-only loss (’OFF’)
-    while  $\{N_{b|S}<N_{S}^{max}\}\;\;\mbox{OR}\;\;\{N_{b}<N^{\rm max}_{b}$}  do
-        $\phi,\mathcal{L}_{\lambda},\mathcal{L}_{S}\leftarrow tuningStep(H_{b},P,net,\phi,\alpha,\beta,full)$
-        $N_{b}\leftarrow N_{b}+1$
-        if {full} then
-           $N_{b|S}\leftarrow N_{b|S}+1$
-           if $\mathcal{L}_{\lambda}/\mathcal{L}_{\lambda}^{EXP}>\gamma_{high}$then
-               $full\leftarrow^{\prime}OFF^{\prime}$triangleright swith-off
-           else if $\mathcal{L}_{\lambda}/\mathcal{L}_{\lambda}^{EXP}<\gamma_{low}$then
-               $full\leftarrow{{}^{\prime}ON^{\prime}}$triangleright swith-on
-           end if
-        end if
-    end while
-    $\hat{H}_{b}\leftarrow predict(H_{b},P,net,\phi)$triangleright inference on band $b$
-end for
-return $\hat{H}$
 TABLE III: Main hyperparameters of the proposed method.
 Param Value Meaning
-$\gamma_{{}_{\rm HIGH}}$ 0.65 Superior limit fixed for ${\cal L}_{\lambda/E}$
-$\gamma_{{}_{\rm LOW}}$ 0.59 Inferior limit for ${\cal L}_{\lambda/E}$
-$\beta_{0}$ 2 Starting $\beta$, weight of the spatial loss term
-$\epsilon$ 0.007 Tolerated growth factor of ${\cal L}_{\lambda}$ for $\beta$ adaptation
-$N_{S}^{\rm max}$ 20 Target # of tuning iterations with active spatial loss
-$N_{0}$ 80 Max # of iterations for bands with perfect correlation with the preceding one
-$\eta$ 30 Overhead factor to globally constrain extra iterations
+$\gamma_{{}_{\rm HIGH}}$ 0.65 Upper limit for ${\cal L}_{\lambda/E}$
+$\gamma_{{}_{\rm LOW}}$ 0.59 Lower limit for ${\cal L}_{\lambda/E}$
+$\beta_{0}$ 2 Initial $\beta$, weight of the spatial loss term
+$\epsilon$ 0.007 Growth factor of ${\cal L}_{\lambda}$ for $\beta$ adaptation
+$N_{S}^{\rm max}$ 20 Target # of tuning iter.s with active spatial loss
+$N_{0}$ 80 Max # of tuning iter.s when $\mathbf{c}_{b}=1$
+$\eta$ 30 Global param. controlling total # of tuning iter.s
 IV-B Assessment Methodology
 To assess the quality of the proposed solution we take advantage of the Benchmarking toolbox [40] which provides a publicly available set of rich and diverse PAN-HS PRISMA images, a large set of SotA comparative methods, and several performance evaluation tools based on the most credited assessment indexes.
-PRISMA was chosen for this evaluation since it provides a real panchromatic image acquired through a dedicated sensor operating in the 400 - 700 nm spectral range.
-In addition to the PAN sensor, PRISMA is equipped with two hyperspectral instruments: a VNIR sensor covering wavelengths from 400 to 1010 nm with 66 bands, and a SWIR sensor spanning 920 to 2505 nm with 173 bands, resulting in a total of 239 spectral channels.
-The dataset released by [42] consists of 16 PRISMA images, 12 designated for training and validation, and 4 for testing.
-To enhance data reliability, noisy spectral bands are removed, reducing the dataset to 159 clean and informative bands for benchmarking purposes.
-Additionally, to further investigate the generalizability of the proposed method, we considered an additional test dataset in which the PAN image was simulated using the RGB bands of the Copernicus Sentinel-2 mission (665,560 and 490 nm with 10m spatial resolution) and subsequently co-registered with the PRISMA HS data.
+TABLE IV: PAN-MS PRISMA benchmark datasets for testing.
+Date Location Land covers
+05/09/22 Cagliari, Italy Roofs, Streets, Crops, Hills, Water
+24/08/23 Udine, Italy Roofs, Streets, Crops, Water
+08/09/23 Ford County, Kansas Streets, Crops
+20/11/23 Macuspana, Tabasco Streets, Crops
+RR test image size: 600$\times$600 (from a 3600$\times$3600 real FR image)
+Real FR test image size: 1200$\times$1200
+Resolution ratio $R=6$
+Bands $B=159/239$
+Resolution: 30 m (RR) or 5 m (FR)
+PRISMA satellites are equipped with a sensor operating in the 400–700 nm spectral range that provides a panchromatic image with 5 m spatial resolution, and two hyperspectral sensors that acquire 66 bands at VNIR wavelengths, 400–1010 nm, and 173 bands at SWIR wavelengths, 920–2505 nm, with 30 m resolution.
+The dataset released by [40] includes 16 images (12 for training and validation, 4 for testing) comprising 159 spectral bands after removing noisy and uninformative channels.
+For our experiments we use the four test images of the Toolbox, see Tab.
+IV.
+From each original PRISMA PAN-HS pair (PAN/HS size: 3600$\times$3600/600$\times$600) two test datasets were created: a FR one (PAN/HS size: 1200$\times$1200/200$\times$200) obtained by cropping; a RR one (PAN/HS/GT size 600$\times$600/100$\times$100/600$\times$600) obtained by resolution downgrading of the whole PAN-HS pair, with the original FR HS playing the role of GT.
 Tab.
-IV gathers the comparative methods which can be roughly grouped in four categories: component substitution (CS), multiresolution analysis (MRA), model-based optimization (MBO), and deep learning (DL).
+V gathers the comparative methods which can be roughly grouped in four categories: component substitution (CS), multiresolution analysis (MRA), model-based optimization (MBO), and deep learning (DL).
 The interested Reader is referred to [40] for details on these methods.
-TABLE IV: HS pansharpening methods.
+TABLE V: HS pansharpening methods.
 Name Ref Summary
@@
 Accordingly, the quality indexes, gathered in Tab.
-V, are split in two groups for reduced resolution (RR) and full resolution (FR) assessment.
+VI, are split in two groups for reduced resolution (RR) and full resolution (FR) assessment.
 As well known, the assessment of spatial quality in the full resolution context (hence, without a GT) is still an open and challenging problem.
@@
 Even so, no numerical index or combination of indexes provides a fully reliable measure of spatial quality, and we will resort to visual inspection of sample results to get further indications on spatial quality.
-TABLE V: HS pansharpening quality assessment indexes.
+TABLE VI: HS pansharpening quality assessment indexes.
 RR assessment
@@
 $D_{\rho}$ Correlation-based spatial distortion index [83]
-TABLE VI: PAN-MS PRISMA benchmark datasets for testing.
-Date Location Land covers
-05/09/22 Cagliari, Italy Roofs, Streets, Crops, Hills, Water
-24/08/23 Udine, Italy Roofs, Streets, Crops, Water
-08/09/23 Ford County, Kansas Streets, Crops
-20/11/23 Macuspana, Tabasco Streets, Crops
-RR test image size: 600$\times$600 (from a 3600$\times$3600 real FR image)
-Real FR test image size: 1200$\times$1200
-Resolution ratio $R=6$
-Bands $B=159/239$
-Resolution: 30 m (RR) or 5 m (FR)
-Experiments are carried out on the test datasets of the Toolbox, summarized in Tab.
-VI, comprising four acquisitions/scenes.
-From each original PRISMA PAN-HS pair (PAN/HS size: 3600$\times$3600/600$\times$600) two test datasets were created: a FR one (PAN/HS size: 1200$\times$1200/200$\times$200) obtained by cropping; a RR one (PAN/HS/GT size 600$\times$600/100$\times$100/600$\times$600) obtained by resolution downgrading of the whole PAN-HS pair, with the original FR HS playing the role of GT.
-Regarding the additional test dataset with the Sentinel-2 simulated PAN, we selected a Sentinel-2 acquisition covering the same area as the PRISMA Kansas dataset (in Tab.
-VI), acquired with a one-day temporal shift, and co-registered it accordingly.
-To preserve the original PRISMA ratio of 6, the hyperspectral data were then downscaled by a factor of 2, resulting in a spatial resolution of 60 meters (PAN/HS size: 600$\times$600/100$\times$100).
 IV-C Reduced-resolution Quality Assessment
@@
 (Ideal)   0 0 0 0 0   0 0 0 0 0   0 0 0 0 0 EXP   0.0126 0.0077 0.0054 0.0106 0.0091   0.1306 0.1273 0.0704 0.1947 0.1308   0.7973 0.8839 0.7963 0.7972 0.8187 GSA   0.0227 0.0134 0.0078 0.0122 0.0140   0.0056 0.0074 0.0077 0.0184 0.0098   0.0593 0.6009 0.2075 0.0791 0.2367 BT-H   0.0633 0.0936 0.0195 0.0288 0.0513   0.0000 0.0000 0.0000 0.0002 0.0001   0.0579 0.0954 0.0544 0.0676 0.0688 BDSD-PC   0.0439 0.0256 0.0220 0.0487 0.0351   0.0024 0.0000 0.0001 0.0003 0.0007   0.1280 0.3878 0.1726 0.1192 0.2019 PRACS   0.0118 0.0155 0.0047 0.0088 0.0102   0.0083 0.0163 0.0114 0.0245 0.0151   0.1775 0.3252 0.3409 0.1903 0.2585 MTF-GLP-FS   0.0101 0.0054 0.0033 0.0065 0.0063   0.0214 0.0277 0.0174 0.0322 0.0247   0.0618 0.6000 0.2123 0.0831 0.2393 MTF-GLP-HPM   0.0147 0.0087 0.0067 0.0093 0.0098   0.0215 0.0284 0.0179 0.0389 0.0267   0.0479 0.1022 0.0438 0.0538 0.0619 MTF-GLP-HPM-R   0.0100 0.0057 0.0033 0.0074 0.0066   0.0217 0.0285 0.0176 0.0374 0.0263   0.0669 0.6050 0.2164 0.0950 0.2458 AWLP   0.0132 0.0097 0.0055 0.0092 0.0094   0.0251 0.0318 0.0206 0.0407 0.0295   0.0695 0.1778 0.0754 0.0933 0.1040 MF   0.0655 0.0392 0.0390 0.0331 0.0442   0.0389 0.0269 0.0250 0.0664 0.0393   0.0476 0.0904 0.0391 0.0585 0.0589 HySURE   0.1126 0.0618 0.0510 0.1158 0.0853   0.0022 0.0011 0.0016 0.0022 0.0018   0.1414 0.2276 0.0990 0.2688 0.1842 SR-D   0.0128 0.0087 0.0055 0.0115 0.0096   0.1316 0.1292 0.0720 0.1981 0.1327   0.8234 0.9061 0.8531 0.8560 0.8596 TV   0.0042 0.0028 0.0021 0.0036 0.0032   0.0586 0.0524 0.0358 0.0813 0.0570   0.2590 0.3294 0.2791 0.2001 0.2669 HyperPNN   0.0415 0.0139 0.0256 0.0248 0.0264   0.0036 0.0063 0.0040 0.0067 0.0051   0.1310 0.5080 0.2419 0.2625 0.2859 HSpeNet   0.0174 0.0084 0.0240 0.0154 0.0163   0.0101 0.0160 0.0141 0.0236 0.0159   0.1439 0.5338 0.2348 0.2558 0.2921 DHP-DARN   0.0832 0.0615 0.0467 0.0513 0.0607   0.0030 0.0053 0.0031 0.0130 0.0061   0.1494 0.4713 0.2106 0.2315 0.2657 DIP-HyperKite   0.0208 0.0125 0.0148 0.0146 0.0157   0.0042 0.0050 0.0043 0.0059 0.0048   0.2952 0.5270 0.2217 0.2765 0.3301 Hyper-DSNet   0.0119 0.0050 0.0082 0.0068 0.0080   0.0200 0.0246 0.0195 0.0346 0.0247   0.1957 0.5336 0.2276 0.3353 0.3230 R-PNN   0.0131 0.0091 0.0063 0.0074 0.0090   0.0168 0.0144 0.0129 0.0341 0.0195   0.0511 0.1366 0.0532 0.0504 0.0728 PCA-Z-PNN   0.0129 0.0041 0.0046 0.0080 0.0074   0.0152 0.0116 0.0088 0.0257 0.0153   0.0959 0.6828 0.7935 0.0792 0.4128 $\rho$-PNN   0.0043 0.0027 0.0018 0.0043 0.0033   0.0215 0.0168 0.0173 0.0380 0.0234   0.0790 0.1302 0.0743 0.0532 0.0842
-TABLE IX: Results at Full Resolution on Sentinel-2 / PRISMA dataset. Best score is in bold green, Top-5 in green, and Worst-5 in red
-.
-Method   $D_{\lambda}$ $D_{S}$ $D_{\rho}$ (Ideal)   0 0 0 EXP   0.0039 0.0985 0.7919 GSA   0.0068 0.0002 0.2833 BT-H   0.0338 0.0000 0.0298 BDSD-PC   0.0095 0.0011 0.1236 PRACS   0.0047 0.0023 0.2629 MTF-GLP-FS   0.0019 0.0134 0.2820 MTF-GLP-HPM   0.0039 0.0140 0.0384 MTF-GLP-HPM-R   0.0019 0.0139 0.2861 AWLP   0.0034 0.0154 0.0626 MF   0.0278 0.0340 0.0439 HySURE   0.0553 0.0068 0.1323 SR-D   0.0043 0.1015 0.8624 TV   0.0018 0.0536 0.2117 HyperPNN   0.0824 0.0027 0.1967 HSpeNet   0.0920 0.0033 0.0799 DHP-DARN   0.1136 0.0035 0.3505 DIP-HyperKite   0.0044 0.0032 0.1657 Hyper-DSNet   0.0284 0.0121 0.1894 R-PNN   0.0032 0.0079 0.0516 PCA-Z-PNN   0.0013 0.0067 0.8200 $\rho$-PNN   0.0013 0.0071 0.0521
 IV-D Full-resolution quality assessment
@@
 In fact, neither of them is without flaws.
-For example, $D_{S}$ quantifies the distance between the PAN image and the optimal linear combination of the pansharpened spectral bands that best approximates it [92].
-This is physically motivated by the fact that a high-resolution hyperspectral (HS) sensor would capture narrowband signals that collectively contribute to the broadband response of the PAN sensor.
-Therefore, an ideal set of sharpened HS bands should be able to reconstruct the PAN image through an appropriate linear combination.
-However, if one band (among hundreds) is an exact copy of the PAN, then $D_{S}=0$, no matter what the other bands look like.
+For example, $D_{S}$ measures the distance between the PAN and the best (closest to the PAN) linear combination of the pansharpened spectral bands [90].
+Therefore, if one band (among hundreds) is an exact copy of the PAN, then $D_{S}=0$, no matter what the other bands look like.
 This happens for several entries in Tab.
@@
 Similar warnings apply to the second indicator.
-In fact, $D_{\rho}$ is defined as the average of the local (calculated on a $\sigma{\times}\sigma$ domain) correlation coefficient between PAN and the pansharpened HS bands [85].
+In fact, $D_{\rho}$ is defined as the average of the local correlation coefficient between PAN and the pansharpened HS bands [83].
 From a physical perspective, this metric evaluates the extent to which a $\sigma{\times}\sigma$ patch from any band of $\widehat{\mathbf{\mathbf{H}}}$ can be linearly predicted from the corresponding patch in the PAN image.
@@
 However, this is not desirable for bands that have no spectral overlap with the PAN and may differ significantly from it.
-Although both indices have their limitations, we believe that their combined use offers a more comprehensive assessment of spatial consistency.
 Keeping this in mind, we comment on the numerical results with some caution and await confirmation from the visual analysis to draw firm conclusions.
@@
 Only R-PNN provides a better $D_{\rho}$ on average, but this is also due to the perfect alignment between the metric and the spatial loss of R-PNN, a property that does not hold with the new spatial loss used in the proposal.
-The numerical results for the Sentinel-2/PRISMA dataset are reported in Table IX.
-In this experiment, it is evident that none of the implemented methods, except for the proposed one, achieves a top-5 ranking in both spectral quality ($D_{\lambda}$) and spatial quality ($D_{S}$ and/or $D_{\rho}$).
 IV-E Visual results
@@
 VIII) plus TV, top-1 $D_{\lambda}$, for a total of 10 methods including the proposed one.
-[^3] The full set of results is available on the web repository at https://github.com/giu-guarino/rho-PNN..
-Figure 12 presents the results for the Sentinel-2/PRISMA dataset.
-As in Figures 10 and 11, we include 10 methods following the same selection criteria, with the exception of replacing TV with PCA-Z-PNN (the second-best performer in terms of $D_{\lambda}$ after the proposed method).
-Together with the images, we also report the spectral quality ($D_{\lambda}$) and spatial quality ($D_{S}/D_{\rho}$) scores computed over all bands of the full scene.
-Additionally, Figure 13 presents the visual results of the top-5 methods in terms of $Q^{2n}$ on the RR Udine dataset, along with their corresponding error maps.
-As in previous cases, we also report the evaluation metrics ($ERGAS$, $SAM$, and $Q^{2n}$) computed over all spectral bands and the full scene.
+[^3] The full set of results is available on the web repository at
+https://github.com/giu-guarino/rho-PNN.
+Together with the images, we also report the spectral ($D_{\lambda}$) and spatial ($D_{S}$ and $D_{\rho}$) quality scores computed over all bands of the full scene.
 PAN HS BT-H BDSD-PC HySURE DIP-HyperKite
@@
 The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm).
-For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab.
+For each result the corresponding scores “($D_{\lambda}$), ($D_{S}$, $D_{\rho}$)” are reported (from Tab.
 VIII).
@@
 The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm).
-For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab.
+For each result the corresponding scores “($D_{\lambda}$), ($D_{S}$, $D_{\rho}$)” are reported (from Tab.
 VIII).
+GT HyperPNN HSpeNet DIP-HyperKite Hyper-DSNet $\rho$-PNN
+[…]
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 12: Pansharpening results on the RR Udine dataset (120$\times$180 close-up) for visible bands (a) and NIR-SWIR bands (b).
+In both cases, ground truth and pansharpening results (top row) with error maps (bottom row).
+For each result the corresponding scores “(ERGAS, SAM, $Q2^{n}$)” are reported.
+From the visual inspection of images we can make the following observations:
+  • Results with low $D_{\rho}$ are generally sharper, resembling closely the spatial structures of the PAN.
+This is true for scenes and in both the visible and NIR-SWIR spectral ranges.
+In particular, this is the case of BT-H, MF, MTF-GLP-HPM, R-PNN and proposed.
+  • Results with low $D_{S}$ reproduce the PAN spatial structures faithfully in the visible spectrum, but seem to fail outside that range.
+This is the case of BDSD-PC, HySURE and DIP-HyperKite.
+BT-H seems an exception, characterized also by a low $D_{\rho}$, however it presents a large spectral distortion.
+  • The top-$D_{\lambda}$ method, TV, whose spectral quality level is nearly the same as the proposed one, does not provide effective sharpening, providing blurred images.
+Based on the above observations, we single out three methods, MTF-GLP-HPM, R-PNN and our proposal, that appear to be both robust and effective, performing uniformly well on both scenes and all displayed spectral bands in terms of both spectral and spatial quality.
+Samples of these methods are shown close to each other but the differences are so small that further examination seems like a futile exercise that could lead to claims related only to the choice of scene/bands/clips.
+On the other hand, all three methods have very similar spatial quality indicators.
+However, we point out that the proposed method has a much smaller average spectral distortion (0.0033, see Tab.VIII) than these competitors (0.0098 and 0.0090, respectively).
+Furthermore, only the proposed method ensures uniform quality across all bands, a feature not captured by the global indicators.
+We complete this analysis by showing, in Fig.
+12, some error maps obtained by subtracting the ground truth (GT) from the pansharpened image.
+In this case, we use the Udine dataset, where some interesting phenomena are visible, and work at reduced resolution, where the original HS component can be used as GT.
+To save space only 5 methods are considered, the proposed one and the best four competitors according to the Q2n metric.
+For bands in the visible range (top) only small errors occur, and the error maps appear almost uniformly gray, with the exception of HyperPNN where some color aberrations are visible.
+As expected, larger errors appear for the NIR–SWIR bands (bottom) that do not overlap the PAN.
+Interestingly, some sharp structures visible in the GT (probably, waterways) are almost completely lost in the fused images, uniformly for all methods.
+On the other hand, such structures are hardly distinguishable in the visible bands and therefore in the PAN, suggesting that the relevant information may be truly missing in this case.
+IV-F Generalization Ability
+As already said, $\rho$-PNN does not require any preliminary training phase but adapts on the fly to the target data.
+Therefore, we expect it to generalize smoothly to new sources, a key requirement of data-driven methods.
+To investigate this issue experimentally we would need an alternative dataset, but we could find none available online, so we built a new one ad hoc, taking advantage of the good coverage and wide availability of Copernicus Sentinel-2 imagery.
+We found a Sentinel-2 image covering the same area as our PRISMA Kansas dataset with a temporal shift of one day.
+We used the RGB (665, 560 and 490 nm) high-resolution bands of the Sentinel-2 image to synthesize a 10-meter resolution PAN, and co-registered it with the PRISMA HS dataset.
+Finally, to preserve the resolution ratio of 6 used in the toolbox [42] for all methods, the hyperspectral data were downscaled by a factor of 2, resulting in a spatial resolution of 60 meters.
+On this heterogeneous dataset we repeated the same experiments carried out on the homogeneous PRISMA datasets working, however, only at full resolution, for lack of sufficient data.
+The Results are reported in Table IX and Figure 13 with the by now usual format, where comparison methods are selected based on the results of Table IX.
+Both numerical and visual results confirm our expectations: $\rho$-PNN keeps working very well and is the only method that achieves top-5 ranking for both spectral and spatial quality.
+On the contrary, several deep learning-based methods appear to struggle with these new data, especially in terms of spectral quality.
+This is especially obvious for HyperPNN, with pansharpened images that are severely corrupted, see Figure 13.
+Model-based methods, on the other hand, are not significantly affected by the data mismatch.
+In extreme synthesis, these results show that the proposed method keeps ensuring a very good and stable performance also on this new dataset, confirming the expected good generalization ability.
+TABLE IX: Results at Full Resolution on Sentinel-2 / PRISMA dataset. Best score is in bold green, Top-5 in green, and Worst-5 in red
+.
+Method   $D_{\lambda}$ $D_{S}$ $D_{\rho}$ (Ideal)   0 0 0 EXP   0.0039 0.0985 0.7919 GSA   0.0068 0.0002 0.2833 BT-H   0.0338 0.0000 0.0298 BDSD-PC   0.0095 0.0011 0.1236 PRACS   0.0047 0.0023 0.2629 MTF-GLP-FS   0.0019 0.0134 0.2820 MTF-GLP-HPM   0.0039 0.0140 0.0384 MTF-GLP-HPM-R   0.0019 0.0139 0.2861 AWLP   0.0034 0.0154 0.0626 MF   0.0278 0.0340 0.0439 HySURE   0.0553 0.0068 0.1323 SR-D   0.0043 0.1015 0.8624 TV   0.0018 0.0536 0.2117 HyperPNN   0.0824 0.0027 0.1967 HSpeNet   0.0920 0.0033 0.0799 DHP-DARN   0.1136 0.0035 0.3505 DIP-HyperKite   0.0044 0.0032 0.1657 Hyper-DSNet   0.0284 0.0121 0.1894 R-PNN   0.0032 0.0079 0.0516 PCA-Z-PNN   0.0013 0.0067 0.8200 $\rho$-PNN   0.0013 0.0071 0.0521
 PAN HS GSA BT-H BDSD-PC PRACS
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 12: Pansharpening results on the Sentinel-2 / PRISMA dataset (240$\times$360 close-up): Simulated PAN image obtained from Sentinel-2 RGB bands (top left) with paired lower resolution PRISMA HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results.
+Figure 13: Pansharpening results on the synthetic Sentinel-2/PRISMA dataset (240$\times$360 close-up): Simulated PAN image obtained from Sentinel-2 RGB bands (top left) with paired reduced resolution PRISMA HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results.
 The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm).
-For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab.
-VIII).
-GT HyperPNN HSpeNet DIP-HyperKite Hyper-DSNet $\rho$-PNN
-[…]
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 13: Pansharpening results on the RR Udine dataset (120$\times$180 close-up) and error maps (Top 5 $Q2^{n}$).
-RGB (a) and False color (b) results are displayed.
-For each result the corresponding scores “ERGAS, SAM, $Q2^{n}$” are reported.
-From the visual inspection of images we can make the following observations:
-  • Results with low $D_{\rho}$ are generally sharper, resembling closely the spatial structures of the PAN.
-This is true for both PRISMA FR scenes and Sentinel-2/PRISMA scene and in both the visible and NIR-SWIR spectral ranges.
-In particular, this is the case of BT-H, MF, MTF-GLP-HPM, R-PNN and proposed.
-  • Results with low $D_{S}$ reproduce the PAN spatial structures faithfully in the visible spectrum, but seem to fail outside that range.
-This is the case of BDSD-PC, HySURE and DIP-HyperKite form PRISMA FR scene, whereas GSA, BDSD-PC, and PRACS are the cases for Sentinel-2/PRISMA scene.
-BT-H seems an exception, characterized also by a low $D_{\rho}$, however it presents a large spectral distortion.
-  • For PRISMA FR scenes, the top-$D_{\lambda}$ method, TV, whose spectral quality level is nearly the same as the proposed one, does not provide effective sharpening, providing blurred images.
-For the RR evaluation framework, we selected image crops from the Udine area (see Fig.
-13), displaying results for the visible bands in the top row and the NIR–SWIR bands in the bottom row.
-The left image shows the ground truth (i.e., the original HS image), followed by the pansharpening results of the five best-performing methods according to the Q2n metric, and finally, the output of our proposed method.
-The visual quality of the results produced by the top-5 algorithms is generally comparable with that of our method, confirming the trends observed in the quantitative results reported in Tab.
-VII.
-However, a closer inspection reveals that our method tends to preserve spectral fidelity more effectively.
-In contrast, the output of HyperPNN exhibits noticeable color artifacts.
-Although some of the other methods provide slightly sharper spatial details, they also introduce visible spectral aberrations, particularly in homogeneous regions such as the sea.
-Based on the above observations, we single out three methods, MTF-GLP-HPM, R-PNN and our proposal, that appear to be both robust and effective, performing uniformly well on all scenes and all displayed spectral bands in terms of both spectral and spatial quality.
-Samples of these methods are shown close to each other but the differences are so small that further examination seems like a futile exercise that could lead to claims related only to the choice of scene/bands/clips.
-On the other hand, all three methods have very similar spatial quality indicators.
-However, we point out that the proposed method has a much smaller average spectral distortion (0.0033, in Tab.VIII and 0.0013 in Tab.IX) than these competitors (0.0098 and 0.0090, respectively, in Tab.VIII, and 0.0039 and 0.0032, respectively in Tab.IX).
-Furthermore, only the proposed method ensures uniform quality across all bands, a feature not captured by the global indicators.
-In summary, the proposed method demonstrates more consistent and robust performance compared to current SotA.
-It effectively preserves spectral fidelity without compromising spatial quality, and unlike other deep learning-based approaches, it does not require any training data for pansharpening inference.
-Furthermore, as confirmed by an additional test involving a PAN image acquired from a different sensor on a different date, the method exhibits a strong degree of generalizability.
-On the downside, under the current hyperparameter configuration, the approach entails a relatively high computational cost compared to both model-based and several deep learning-based techniques.
-However, by relaxing the spectral quality constraints (via $\gamma_{\text{high}}$ and $\gamma_{\text{low}}$), it is possible to accelerate the algorithm significantly.
+For each result the corresponding scores “($D_{\lambda}$), ($D_{S}$, $D_{\rho}$)” are reported (from Tab.
+IX).
+IV-G Discussion
+This large body of experimental results fully confirms our expectations about the proposed method.
+For the sake of clarity, we summarize here its pros and cons, also with reference to competing methods.
+From a structural point of view, $\rho$-PNN presents several qualifying points, some of which are shared only with its ancestor R-PNN, and with PCA-Z-PNN:
+  • it runs from scratch, without the need for prior training: this is obtained by means of a longer tuning run on the first band that affects only marginally the overall computational cost due to the very large number of bands;
+  • it works with an arbitrary number of bands: as the algorithm works band-wise with no joint processing of the HS stack, the number of spectral bands does not need to be fixed to build the model;
+  • it operates at full resolution with no resolution downgrading or simulation operations: as the proposed loss is fully unsupervised and based of input-output self-consistency, traditional resolution downgrade protocols (e.g., [82]) to give rise to labeled, but synthetic, dataset are avoided.
+Furthermore, in terms of performance
+  • it is on par with the best SotA methods, both model-based and data-driven;
+  • it guarantees a uniformly high spectral quality at all wavelengths, as only the TV method does;
+  • it shows excellent generalization ability.
+On the down side, $\rho$-PNN has a relatively high computational cost compared to model-based techniques and even several data-driven techniques.
+This cost can be significantly reduced, however, by slightly relaxing the spectral quality constraints.
 V Conclusions
@@
 From this base, we improve with two key innovations: a new loss, better adapted to the characteristics of HS images, that takes into account the correlation inversion phenomena that often occur in bands outside the visible range; and a hysteresis-based tuning scheme where the balance of spatial and spectral loss components changes dynamically to overcome unwanted low-quality solutions.
-Advantages
-  • It works from scratch (Only PCA-Z-PNN does the same);
-  • It works with an arbitrary number of bands (Only PCA-Z-PNN and R-PNN do the same);
-  • It ensures a good generalization ability across diverse sensors;
-  • It works at full resolution: No downgrade nor simulation operations are needed (same as R-PNN and PCA-Z-PNN);
-  • It ensrures a controllable and uniform level of spectral control along all wavelengths.
-Drawbacks
-  • It is slower than pre-trained methods at inference time;
 Extensive experimental assessment over a SotA benchmarking toolbox proves the proposed method to ensure excellent results both in terms of visual quality and according to the most credited quality indicators.